### v2

# Hadamard-type formulas for real eigenvalues of canonically symplectic operators

Abstract.
We give first-order asymptotic expansions for the resolvent and Hadamard-type formulas for the eigenvalue curves of one-parameter families of canonically symplectic operators. We allow for parameter dependence in the boundary conditions, bounded perturbations and trace operators associated with each off-diagonal operator, and give formulas for derivatives of eigenvalue curves emanating from the discrete eigenvalue of the unperturbed operator in terms of Maslov crossing forms. We derive the Hadamard-type formulas using two different methods: via a symplectic resolvent difference formula and asymptotic expansions of the resolvent, and using Lyapunov-Schmidt reduction and the implicit function theorem. The latter approach facilitates derivative formulas when the eigenvalue curves are viewed as functions of the spectral parameter. We apply our abstract results to derive a spectral index theorem for the linearised operator associated with a standing wave in the nonlinear Schrödinger equation on a compact star graph.
Key words and phrases:
Eigenvalues, Maslov index, crossing forms, star graphs
$S.$ was supported in part by NSF grants DMS–2510344, DMS–2418900, Simons Foundation grant MP–TSM–00002897, grant no. $2024154$ from the U.S.-Israel Binational Science Foundation Jerusalem, Israel, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program grant.
1. Introduction
Recently many works have used the Maslov index to study nonlinear PDEs. In this paper, we are motivated by two recent works [CCLM23, LS20first]. The first discusses Hadamard-type formulas for rates of change of eigenvalues of self-adjoint realizations of elliptic operators on bounded domains and abstract symmetric operators in Hilbert spaces [LS20first]; the second discusses similar formulas for the rates of change of real eigenvalues, with respect to perturbations of the domain, of a canonically symplectic operator [KKS04] arising in the stability analysis of standing waves in the NLS equation on a compact interval subject to Dirichlet boundary conditions [CCLM23]. Here we extend the results of the latter by using the framework of the former: we study the variation of real eigenvalues of canonically symplectic operators where the off-diagonal blocks are self-adjoint extensions of densely-defined symmetric operators. Our results bridge the celebrated classical Rayleigh-Hadamard-Rellich theory for eigenvalue variation [Rayleigh45, hadamard1908, Rellich69] with the Arnold-Maslov-Keller index theory from symplectic geometry [Maslov65, A67, Keller58] for operators with the canonical symplectic structure. Since such operators are not self-adjoint, most of the monotonicity properties present in the self-adjoint setting are lost, leading to far richer spectral behaviour. Furthermore, as shown in [CCLM23] these local formulas can provide crucial information in determining global counts of eigenvalues. Before discussing our abstract setting and main results of the paper, we provide some examples, accompanied with numerical experiments, to motivate our analysis.
1.1. Motivating examples
Let $\mathcal{G}$ be a star graph with $m$ edges, where each edge is assigned a positive length and a direction. Edge $i$ has length $\ell_{i}\in(0,\infty)$ and is represented by the interval $0,\ell_{i}]$, where 0 corresponds to the central vertex $\ell_{i}$ the free vertex. The Sobolev spaces of functions on $\mathcal{G}$ are denoted by
$L^{2}(\mathcal{G})\coloneqq\bigoplus_{i=1}^{m}L^{2}(0,\ell_{i}),\qquad\widehat{H}^{k}(\mathcal{G})\coloneqq\bigoplus_{i=1}^{m}H^{k}(0,\ell_{i}),\ k\in\mathbb{N},$
where $H^{k}(0,\ell_{i})$ is the standard $L^{2}$ based Sobolev space of order $k\in\mathbb{N}$ on the interval $(0,\ell_{i})$. The boundary $\partial\mathcal{G}$ is defined by
$\partial\mathcal{G}\coloneqq\cup_{1\leq i\leq m}\{a_{i},b_{i}\},$ (1.1)
where $a_{i},b_{i}$ denote the end points of the $i$th edge. It is convenient to treat the $2m$ dimensional vectors representing the values of the functions at the endpoints of each edge as functions on the boundary $\partial\mathcal{G}$, in particular, $L^{2}(\partial\mathcal{G})\cong\mathbb{C}^{2m}.$
On $\mathcal{G}$ we consider the nonlinear Schrödinger equation with power nonlinearity:
$i\Psi_{t}=\Psi_{xx}+|\Psi|^{2p}\Psi,\qquad x\in\mathcal{G},\quad t\in\mathbb{R},\quad p>0,$ (1.2)
where $\Psi(x,t)=\left(\Psi_{1}(x,t),\dots,\Psi_{m}(x,t)\right)^{\top}\in\mathbb{C}^{m}$, $|\Psi|^{2p}\Psi\coloneqq\left(|\Psi_{1}|^{2p}\Psi_{1},\dots,|\Psi_{m}|^{2p}\Psi_{m}\right)^{\top}$. A standing wave solution is given by
$\Phi(x,t)=e^{i\beta t}\phi(x),\qquad\beta\in\mathbb{R},$ (1.3)
where the wave profile $\phi(x)\coloneqq(\phi_{1}(x),\dots,\phi_{m}(x))^{\top}\in\mathbb{R}^{m}$ solves the standing wave equation (defined edge-wise)
$\phi^{\prime\prime}+\beta\phi+\phi^{2p+1}=0,\qquad\phi^{2p+1}\coloneqq(\phi_{1}^{2p+1},\dots,\phi_{m}^{2p+1})^{\top}.$ (1.4)
We let $\Phi$ satisfy Dirichlet conditions at the free vertices and $\delta$-type conditions at the central vertex,
$\begin{cases}&\phi_{1}(\ell_{1})=\dots=\phi_{m}(\ell_{m})=0,\\ &\phi_{1}(0)=\dots=\phi_{m}(0),\\ &\sum_{k=1}^{m}\phi_{k}^{\prime}(0)=\alpha\phi_{1}(0),\qquad\alpha\in\mathbb{R},\end{cases}$ (1.5)
and note that when $\alpha=0$ one has Neumann-Kirchhoff conditions corresponding to current conservation at the central vertex.
Assuming a solution to (1.4) – (1.5) exists111see Remark 5.1, linearising (1.2) about such a solution using a complex-valued perturbation that satisfies that same vertex conditions leads to the eigenvalue problem
$(\mathcal{N}+V)\mathbf{u}=\lambda\mathbf{u},\qquad\mathbf{u}=(u,v)^{\top}\in\operatorname{dom}\mathcal{N},$ (1.6)
where $\mathcal{N}$ and $\mathcal{V}$, acting in $L^{2}(\mathcal{G})\oplus L^{2}(\mathcal{G})$, are defined as follows
$\mathcal{N}$ $=\begin{pmatrix}0&-\mathcal{A}\\ \mathcal{A}&0\end{pmatrix},\quad\begin{cases}\mathcal{A}=-\partial_{xx},\quad\operatorname{dom}\mathcal{A}=\left\{u\in\widehat{H}^{2}(\mathcal{G}):\operatorname{tr}u\in\mathcal{L}\right\},\\ \operatorname{dom}\mathcal{N}=\operatorname{dom}\mathcal{A}\times\operatorname{dom}\mathcal{A}=\{\mathbf{u}\in\big(\widehat{H}^{2}(\mathcal{G})\big)^{2}:\operatorname{T}\mathbf{u}\in\mathcal{L}\oplus\mathcal{L}\},\end{cases}$ (1.7a)
$V$ $=\begin{pmatrix}0&-F\\ G&0\end{pmatrix},\qquad\begin{cases}F=-\phi(x)^{2p}-\beta,\\ G=-(2p+1)\phi(x)^{2p}-\beta.\end{cases}$ (1.7b)
In (1.7a) $\operatorname{tr}$ is a trace operator on $\widehat{H}^{2}(\mathcal{G})$ that maps a function to the value of the function and its first derivative on the boundary $\partial\mathcal{G}$,
$\operatorname{tr}:\widehat{H}^{2}(\mathcal{G})\to L^{2}(\partial\mathcal{G})\oplus L^{2}(\partial\mathcal{G}),\quad\operatorname{tr}u=(\Gamma_{0}u,\Gamma_{1}u)^{\top},$ (1.8)
where, for $u=(u_{1},\dots,u_{m})^{\top}\in\widehat{H}^{2}(\mathcal{G})$,
$\begin{split}\Gamma_{0}&:\widehat{H}^{2}(\mathcal{G})\to L^{2}(\partial\mathcal{G}),\qquad\Gamma_{0}u=(u_{1}(0),\dots,u_{m}(0),u_{1}(\ell_{1}),\dots,u_{m}(\ell_{m}))^{\top},\\ \Gamma_{1}&:\widehat{H}^{2}(\mathcal{G})\to L^{2}(\partial\mathcal{G}),\qquad\Gamma_{1}u=(u_{1}^{\prime}(0),\dots,u_{m}^{\prime}(0),-u_{1}^{\prime}(\ell_{1}),\dots,-u_{m}^{\prime}(\ell_{m}))^{\top}.\end{split}$ (1.9)
The operator $\operatorname{T}$ in (1.7a) is then the following trace operator on $\big(\widehat{H}^{2}(\mathcal{G})\big)^{2}$,
$\operatorname{T}:\big(\widehat{H}^{2}(\mathcal{G})\big)^{2}\to\big(L^{2}(\partial\mathcal{G})\big)^{4},\quad\operatorname{T}\mathbf{u}\coloneqq\operatorname{tr}u\oplus\operatorname{tr}v=(\Gamma_{0}u,\Gamma_{1}u,\Gamma_{0}v,\Gamma_{1}v)^{\top},\quad\mathbf{u}=\begin{pmatrix}u\\ v\end{pmatrix}\in\big(\widehat{H}^{2}(\mathcal{G})\big)^{2},$
Finally, $\mathcal{L}$ is a Lagrangian subspace of the symplectic boundary space $\mathbb{C}^{2m}\times\mathbb{C}^{2m}$ corresponding to the vertex conditions described by (1.5).
The operator $\mathcal{N}$ is said to have the canonical symplectic structure [KKS04]. Its spectrum, consisting solely of eigenvalues due to the compactness of the domain, thus has four-fold symmetry in the complex plane. To determine the spectral stability of the standing wave, one must therefore determine the existence of any eigenvalues lying off the imaginary axis. While classical results [GSS87, Grill88, GSS90] are relevant in the present context, see also [Pel05, pelinovsky, KKS04, KKS05, KP12art], here we opt for a symplectic approach, which is related to the study of Hadamard-type variational formulas for parameter-dependent eigenvalues in an effort to further recent studies of the latter [CCLM23, LS20first]. Namely, we consider the problem on a family of shrinking subdomains, parameterised by $t\in(0,1]$, and study the spectrum of the problem on the full domain when $t=1$ by analysing the $t$-dependent spectrum when $\lambda=0$ and $t\in(0,1]$. This approach has been used to study the spectra of ordinary and partial differential operators in many works, for example, [Smale65, DJ11, JLM13, CJLS14, LSHad17].
To that end, let $\mathcal{G}_{t}$, $t\in(0,1]$, be the star graph with edges $0,t\ell_{i}]$. Restricting the eigenvalue problem for the operator (1.6), (1.7a), (1.7b) to $\mathcal{G}_{t}$ and then rescaling back to $\mathcal{G}$ yields the following $t$-dependent family of eigenvalue problems on $\mathcal{G}$,
$(\mathcal{N}_{t}+V_{t})\mathbf{u}=t^{2}\lambda\mathbf{u},\qquad\mathbf{u}\in\operatorname{dom}\mathcal{N}_{t},$ (1.10)
where
$\mathcal{N}_{t}$ $=\begin{pmatrix}0&-\mathcal{A}_{t}\\ \mathcal{A}_{t}&0\end{pmatrix},\quad\begin{cases}\mathcal{A}_{t}=-\partial_{xx},\quad\operatorname{dom}\mathcal{A}_{t}=\left\{u\in\widehat{H}^{2}(\mathcal{G}):\operatorname{tr}_{t}u\in\mathcal{L}\right\},\\ \operatorname{dom}\mathcal{N}=\operatorname{dom}\mathcal{A}\times\operatorname{dom}\mathcal{A}=\{\mathbf{u}\in\big(\widehat{H}^{2}(\mathcal{G})\big)^{2}:\operatorname{T}_{t}\mathbf{u}\in\mathcal{L}\oplus\mathcal{L}\},\end{cases}$ (1.11a)
$V_{t}$ $=\begin{pmatrix}0&-F_{t}\\ G_{t}&0\end{pmatrix},\qquad\begin{cases}F_{t}=-t^{2}\phi(tx)^{2p}-t^{2}\beta,\\ G_{t}=-t^{2}(2p+1)\phi(tx)^{2p}-t^{2}\beta,\end{cases}$ (1.11b)
and
$\operatorname{tr}_{t}:$ $\widehat{H}^{2}(\mathcal{G})\to L^{2}(\partial\mathcal{G})\times L^{2}(\partial\mathcal{G}),\quad$ $\operatorname{tr}_{t}u\coloneqq\left(\Gamma_{0}u,\frac{1}{t}\Gamma_{1}u\right)^{\top},$ (1.12a)
$\operatorname{T}_{t}:$ $\big(\widehat{H}^{2}(\mathcal{G})\big)^{2}\to\big(L^{2}(\partial\mathcal{G})\big)^{4},\quad$ $\operatorname{T}_{t}\mathbf{u}\coloneqq\operatorname{tr}_{t}u\oplus\operatorname{tr}_{t}u_{2}=\left(\Gamma_{0}u_{1},\frac{1}{t}\Gamma_{1}u_{1},\Gamma_{0}u_{2},\frac{1}{t}\Gamma_{1}u_{2}\right)^{\top}.$ (1.12b)
Before stating our main result relating to this example, we note that $\phi\in\ker(\mathcal{A}+F)$, since by (1.4) we have
$(\mathcal{A}+F)\phi=\phi^{\prime\prime}+\phi^{2p+1}+\beta\phi=0,$ (1.13)
and because $\phi$ satisfies (1.5) we have $\operatorname{tr}\phi=(\Gamma_{0}\phi,\Gamma_{1}\phi)^{\top}\in\mathcal{L}$, hence $\phi\in\operatorname{dom}(\mathcal{A})$. We make the assumption that zero is a simple eigenvalue of $\mathcal{A}+F$; in the case of Neumann-Kirchhoff conditions (where $\alpha=0$), this assumption is generic with respect to the set of edge lengths $\{\ell_{i}\}$ [berk, Theorem 3.1.7]. We also assume invertibility of $\mathcal{A}+G$.
Hypothesis 1.1.
We assume that $\ker(\mathcal{A}+F)=\operatorname{span}\{\phi\}$ and $\ker(\mathcal{A}+G)=\{0\}$.
It follows from the previous assumption that zero is a simple eigenvalue of $\mathcal{N}+V$ with eigenfunction $\boldsymbol{\phi}\coloneqq(0,\phi)^{\top}$. The real spectral index can then be deduced from an analysis of the eigenvalue curves. The following theorem extends the main result of [CCLM23] from the setting of a compact interval with Dirichlet boundary conditions to star graphs with $\delta$-type conditions on the central vertex and Dirichlet conditions at the free vertices.
Theorem 1.2.
Let $\mathcal{N}_{t}+V_{t}$ be the operator defined by (1.11) for $t\in(0,1]$, associated with the standing wave solution (1.3)–(1.5) to (1.2). (When $t=1$, we drop the subscript, i.e. $\mathcal{N}\coloneqq\mathcal{N}_{1},V\coloneqq V_{1}$.) Under Hypothesis 1.1, there exists a $C^{2}$ curve $t(\lambda)$, defined for $0<|\lambda|\ll 1$, satisfying
$\lambda\in\operatorname{Spec}\left(\mathcal{N}_{t(\lambda)}+V_{t(\lambda)}\right),$ (1.14)
and such that $t(0)=1$ and $t^{\prime}(0)=0$. Furthermore, if the coupling constant $\alpha\geq 0$, cf. (1.5), and $t^{\prime\prime}(0)\neq 0$, then the number of positive real eigenvalues of $\mathcal{N}$ is given by
$n_{+}(\mathcal{N})\geq|p_{c}-q_{c}-\mathfrak{c}|,$ (1.15)
where $p_{c}$ and $q_{c}$ are counts of conjugate points defined as
$\hskip 30.00005ptp_{c}\coloneqq\sum_{t_{0}\in(0,1)}\dim\ker(\mathcal{A}+G_{t_{0}}),\qquad q_{c}\coloneqq\sum_{t_{0}\in(0,1)}\dim\ker(\mathcal{A}+F_{t_{0}}),$ (1.16)
and $\mathfrak{c}$ is determined by the concavity of $t(\lambda)$ at $\lambda=0$ as follows,
$\mathfrak{c}=\begin{cases}+1&t^{\prime\prime}(0)<0,\\ 0&t^{\prime\prime}(0)>0.\end{cases}$ (1.17)
Remark 1.3.
The nondegeneracy condition $t^{\prime\prime}(0)\neq 0$ is generic with respect to the parameters $\beta,p,\phi_{1}(0),\phi_{i}^{\prime}(0)$, $i=1,\dots,m-1$ in (1.4)–(1.5). (If $t^{\prime\prime}(0)=0$ then higher derivatives are needed to determine $\mathfrak{c}$ in (1.17).) Regarding (1.16), the counts $p_{c}$ and $q_{c}$ of conjugate points may alternatively be characterised in terms of nontrivial intersections of Lagrangian planes, i.e.
$p_{c}=\{t\in(0,1):\mathbb{K}_{0,t}\cap(\{0\}\times\mathcal{L})\neq\{0\}\},\quad q_{c}=\{t\in(0,1):\mathbb{K}_{0,t}\cap(\mathcal{L}\times\{0\})\neq\{0\}\},$ (1.18)
where $\mathbb{K}_{\lambda,t}$ is the $t$- and $\lambda$-dependent Cauchy data plane (for a precise definition see Section 4).
Remark 1.4.
We note that the $t$-dependent objects in (1.10)–(1.12b) are well-defined for $t>1$ (which corresponds to stretching the graph $\mathcal{G}$), hence we can extend the allowable $t$ values in Theorem 1.2 to $(0,1+\varepsilon]$ for $\varepsilon>0$ small. Thus the eigenvalue curves are indeed well-defined for $1<t\leq 1+\varepsilon$ when $t^{\prime\prime}(0)>0$.
Remark 1.5.
A corollary of Theorem 1.2 is the following Vakhitov-Kolokolov-type (VK) criterion [VK73, pelinovsky], which furnishes a convenient numerical tool – indeed one need only local data at $(\lambda,t)=(0,1)$ – to establish the existence of a positive real eigenvalue. Suppose $p_{c}=1$ and $q_{c}=0$. Let $t(\lambda)$ be the eigenvalue curve, given in Theorem 1.2, which passes through $(\lambda,t)=(0,1)$ and is defined for $0<|\lambda|\ll 1$. Suppose that $t^{\prime\prime}(\lambda)|_{\lambda=0}\neq 0$. Then $\mathcal{N}+V$ has a positive real eigenvalue if $t^{\prime\prime}(0)>0$, while $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ if $t^{\prime\prime}(0)<0$. The first part of this assertion, that is, existence of unstable eigenvalues follows from Theorem 1.2. While the second part can be shown as in [CCLM23, pelinovsky].
(c) $b=5$
(d) $b=3$
(e) $b=1$
Figure 1. (a) Phase curves (in bold) in the phase plane for (1.4) describing a standing wave solution $\phi=(\phi_{1},\phi_{2},\phi_{3})^{\top}$ on a 3-star, where $\phi_{1}(0)=\phi_{2}(0)=\phi_{3}(0)=1$ and $\phi_{1}^{\prime}(0)=-a-b,\phi_{2}^{\prime}(0)=-a-3,\phi_{3}^{\prime}(0)=-\phi_{1}^{\prime}(0)-\phi_{2}^{\prime}(0)=2a+3+b$, with $a\approx 0.8660$; (b) a schematic of the standing wave. In (c), (d) and (e) we give the eigenvalue curves for the standing wave solution described in (a) with $b$ as indicated. The standing waves all have one bump and two tails, and are non-negative on $\mathcal{G}$.
In Fig. 1 we give some numerical examples that highlight Remark 1.5. The figures showcase the real eigenvalue curves in the $\lambda t$-plane, representing the pairs $(\lambda,t)$ such that there exists a nontrivial solution to (1.10), i.e. so that $t^{2}\lambda$ is an eigenvalue of $\mathcal{N}_{t}+V_{t}$, for three standing wave solutions to the power NLS equation (1.2) with $p=3$ and $\beta=-1$ on the star graph with $m=3$ edges. The standing waves satisfy the following conditions at the central vertex:
$\begin{cases}\phi_{1}(0)=\phi_{2}(0)=\phi_{3}(0)=1,\\ \phi_{1}^{\prime}(0)=-a-b,\,\,\phi_{2}^{\prime}(0)=-a-3,\,\,\phi_{3}^{\prime}(0)=-\phi_{1}^{\prime}(0)-\phi_{2}^{\prime}(0)=2a+3+b,\end{cases}$ (1.19)
where $a\approx 0.8660$ and $b=5,3,1$ in Fig. 1 (c), (d) and (e) respectively. These initial conditions are integrated forward on each edge under (1.4) until the first instance at which the functions are zero. Thus the waves satisfy (1.5) with $\alpha=0$ (i.e. Neumann-Kirchhoff conditions), are non-negative on $\mathcal{G}$ and all have one bump (non-monotonic profile) and two tails (monotonic profile) [PavaGoloshch18, Kairzhan19, DKP22]. Integrating along the orbits in the phase plane associated with (1.4) shows that in all cases we have $\ell_{2}\approx 0.262628$, while
$b=5\implies\ell_{1}\approx 0.171588,\,\,\,\ell_{3}\approx 0.399533,$
$b=3\implies\ell_{1}=\ell_{2},\,\,\,\ell_{3}\approx 0.467768,$
$b=1\implies\ell_{1}\approx 0.575249,\,\,\,\ell_{3}\approx 0.573694,$
in (c), (d) and (e) respectively. The phase curves corresponding to the solutions on each edge, as well as a schematic of the standing wave, are given in (a) and (b), respectively, in Fig. 1.
Using a homotopy argument and [berk, Theorem 5.2.8], one can show that $q_{c}=n_{-}(\mathcal{A}+F)=0$. Noting that the number of intersections of the eigenvalue curves with the line $\lambda=0$ for $t<1$ is the quantity $p_{c}+q_{c}$, Fig. 1 (c), (d) and (e) therefore numerically verify that the three different standing waves described above all satisfy the condition of Remark 1.5. If the eigenvalue curves are continuous and there are no points of horizontal tangency for $t<1$ away from $\lambda=0$, it follows that the existence of a positive real eigenvalue on the full domain (given by intersections with $t=1$) may be predicted from the concavity of the eigenvalue curve through the point $(\lambda,t)=(0,1)$: if $t^{\prime\prime}(\lambda)|_{\lambda=0}>0$ then $\mathcal{N}+V$ has a positive real eigenvalue, while if $t^{\prime\prime}(\lambda)|_{\lambda=0}<0$ then $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ [^2] It can be proven that there are no complex eigenvalues under the assumptions of Remark 1.5, see e.g. [CCLM23, pelinovsky]. Figure 1 therefore reflects a change in stability of the underlying wave as $b$ decreases (and the edge lengths $\ell_{1}$ and $\ell_{3}$ increase): the concavity of the eigenvalue curve changes from negative to positive, indicating that a pair of imaginary eigenvalues have bifurcated onto the real axis.
Following the analysis in [LS20first], in this paper we also concern ourselves with variations in the boundary conditions and their effect on the spectrum of canonically symplectic operators. To this end, in Fig. 2 we numerically compute the eigenvalue curves for three different operators $\mathcal{N}_{t}+V_{t}$ given by (1.11), but where the domain is a compact interval, i.e. $\operatorname{dom}A=H^{2}_{0}(0,\ell)$ and $\operatorname{dom}\mathcal{A}_{t}=\left\{u\in H^{2}:\operatorname{tr}_{t}u\in\mathcal{L}_{t}\right\}$, and:
  1. (1)
    $F_{t}=G_{t}=0$, $\ell=\pi$ and $\mathcal{L}_{t}$ is the Lagrangian plane corresponding to Dirichlet conditions at $x=0$ for all $t\in[0,2]$, and $t$-dependent boundary conditions at $x=\ell$ described by $\operatorname{span}\left\{\begin{bmatrix}\sin\left(\frac{\pi}{2}t\right)\vskip 2.84526pt\\ -\cos\left(\frac{\pi}{2}t\right)\end{bmatrix}\right\}\subset\mathbb{R}^{2}$; thus $\mathcal{L}_{t}$ corresponds to Dirichlet conditions at $x=0,\pi$ when $t=0,2$ and Dirichlet at $x=0$ and Neumann at $x=\pi$ when $t=1$;
  2. (2)
    $F_{t}$ and $G_{t}$ are as in (1.11b) with $p=3$ and $\beta=-2$, where $\phi$ is a $T$-periodic non-negative solution to (1.4) satisfying Dirichlet boundary conditions and $\ell=T/2\approx 3.28418$, and $\mathcal{L}_{t}$ is the Lagrangian plane described in item (1) (with right endpoint $x\approx 3.28418$);
  3. (3)
    The same as in item (2) but with $p=1$, $\ell=T/2\approx 1.09868$ and $\mathcal{L}_{t}$ now the Lagrangian plane with $t$-dependent boundary conditions at $x=0$ and $x=\ell$ both described by $\operatorname{span}\left\{\begin{bmatrix}\sin\left(\frac{\pi}{2}t\right)\vskip 2.84526pt\\ -\cos\left(\frac{\pi}{2}t\right)\end{bmatrix}\right\}\subset\mathbb{R}^{2}$; thus $\mathcal{L}_{t}$ corresponds to Dirichlet conditions at $x=0,\pi$ when $t=0,2$ and Neumann conditions at $x=0,\pi$ when $t=1$.
In contrast to the self-adjoint case, the eigenvalue curves are not monotone, and it is therefore much more difficult to predict how many positive real eigenvalues exist when, for example, $t=1$, given the number of positive real eigenvalues when $t=0$. Moreover, such predictions become even more difficult given the following behaviour observed in all numerical examples: as the Lagrangian plane $\mathcal{L}_{t}$ approaches the Dirichlet plane as $t\rightarrow 0^{+}$ or as $t\rightarrow 2^{-}$, a pair (possibly more) of real eigenvalues diverge to $\pm\infty$. In any case, in this paper it is our goal to understand such eigenvalue curves locally by computing expressions for their derivatives, by allowing for $t$-dependence in the boundary conditions in our abstract setting.
Figure 2. Eigenvalue curves for $\mathcal{N}_{t}+V_{t}$ where (a), (b) and (c) correspond to items (1), (2) and (3) described in the text.
1.2. Discussion of abstract setting and main results
Motivated by the above examples and following the analyses in [LS20first, CCLM23], in this paper we derive Hadamard formulas for the real eigenvalue curves of canonically symplectic operators, i.e. off-diagonal block operators where the blocks are self-adjoint extensions of densely-defined symmetric operators with finite deficiency indices. Here, as in [LS20first], we allow for $t$-dependence in the bounded perturbation of the operator, in the trace operator, and in the Lagrangian planes describing the boundary conditions. We adopt two different approaches: on the one hand, following [LS20first] we derive a symplectic resolvent difference formula which hinges on a modified Green’s identity for the operators of interest. We use these formulas to obtain a first order asymptotic expansion of the resolvent, which, in turn, we use to derive an Hadamard formula for the eigenvalue curve $\lambda(t)$. On the other hand, following [CCLM23] we use Lyaponuv-Schmidt reduction and the implicit function theorem to arrive at the same Hadamard-type formula for $\lambda(t)$, and to derive an additional formula for $t(\lambda)$. As outlined in the first example in the previous section, such a formula is necessary in order to determine global counts of eigenvalues. On a conceptual level, the main difference from the analysis in [CCLM23] is that we now allow for $t$-dependence in the Lagrangian plane describing the boundary conditions (where such Lagrangian plane was the (fixed) Dirichlet plane in [CCLM23]).
Let us succinctly describe the abstract setting and highlight the corresponding results. First, we fix a densely defined symmetric operator $A$ acting in a Hilbert space $\mathcal{H}$ (e.g., $A=-\partial_{xx}$, $\mathcal{H}=L^{2}(\mathcal{G})$ as above) and consider a $2\times 2$ block operator matrix acting in $\mathcal{H}^{2}\coloneqq\mathcal{H}\oplus\mathcal{H}$,
$N:\operatorname{dom}(N)=\operatorname{dom}(A)\times\operatorname{dom}(A)\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad N=\begin{pmatrix}0&-A\\ A&0\end{pmatrix}$. (1.20)
Given a trace map $\operatorname{tr}:\operatorname{dom}(A^{*})\rightarrow\mathfrak{H}\times\mathfrak{H}$ corresponding to $A$, (e.g., (1.9)) we first derive, see Proposition 2.2, a version of the Green’s identity for the non-selfadjoint operator $N$,
$\left\langle(N^{*}-\lambda)\tau\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle\mathbf{u},(N^{*}-\lambda)\tau\mathbf{v}\right\rangle_{\mathcal{H}^{2}}=\Omega(\operatorname{T}\mathbf{u},\operatorname{T}\mathbf{v}),\ \lambda\in\mathbb{R},$ (1.21)
where $\Omega=\omega\oplus(-\omega)$ with $\omega$ being the standard symplectic form associated with $A$, see (2.5), $T:=\operatorname{tr}\oplus\operatorname{tr}$, and $\tau:\mathcal{H}^{2}\rightarrow\mathcal{H}^{2}$ is the involution operator defined by $\tau(u,v):=(v,u)$, see (2.14). Using this version of Green’s identity we first obtain a new resolvent difference formula for a pair of extensions of $N$,
$R_{1}(\zeta)-R_{2}(\zeta)=\tau\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\Big((P_{1}JP_{2})\oplus\left(-Q_{1}JQ_{2}\right)\Big)\operatorname{T}R_{2}(\zeta),$ (1.22)
where for a pair $\mathcal{A}_{i},\mathcal{B}_{i}$ of self-adjoint extensions of $A$ and a pair of orthogonal projections $P_{i},Q_{i}\in\mathcal{B}(\mathfrak{H}^{2},\mathfrak{H}^{2})$ related by $\operatorname{tr}(\operatorname{dom}(\mathcal{A}_{i}))=\operatorname{ran}P_{i}$, $\operatorname{tr}(\operatorname{dom}(\mathcal{B}_{i}))=\operatorname{ran}Q_{i}$, we denote
$R_{i}(\zeta):=(\mathcal{N}_{i}-\zeta)^{-1},\ \ \zeta\not\in\operatorname{Spec}(\mathcal{N}_{i}),\ \ \mathcal{N}_{i}:=\begin{pmatrix}0&-\mathcal{B}_{i}\\ \mathcal{A}_{i}&0\end{pmatrix},\ i=1,2.$ (1.23)
Our second main result leverages (1.22) to provide a first-order asymptotic expansion for a one-parameter family of resolvents $t\mapsto R_{t}(\zeta):=(\mathcal{N}_{t}+V_{t}-\zeta)^{-1}$, where $\mathcal{N}_{t}$ is now determined by one-parameter families of self-adjoint extensions $\mathcal{A}_{t},\mathcal{B}_{t}$ of $A$, trace maps $\operatorname{tr}_{t}$, orthogonal projections $P_{t},Q_{t}$ all related via $\operatorname{tr}_{t}(\operatorname{dom}(\mathcal{A}_{t}))=\operatorname{ran}P_{t}$, $\operatorname{tr}_{t}(\operatorname{dom}(\mathcal{B}_{t}))=\operatorname{ran}Q_{t}$, and the bounded perturbation
$\qquad V_{t}=\begin{pmatrix}0&-G_{t}\\ F_{t}&0\end{pmatrix}$, (1.24)
where $F_{t},G_{t}$ are bounded self-adjoint operators in $\mathcal{H}$ (cf., e.g., (1.11b)). Under natural assumptions on the one-parameter families listed above, see Theorem 3.5 for exact statement, we obtain the following formula
$\begin{split}R_{t}(\zeta)&\underset{t\to t_{0}}{=}R_{t_{0}}(\zeta)+\Big(-R_{t_{0}}(\zeta)\dot{V}_{t_{0}}R_{t_{0}}(\zeta)+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)\\ &\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}R_{t_{0}}(\zeta)\Big)(t-t_{0})+o(t-t_{0})\quad\text{in}\,\,\mathcal{B}(\mathcal{H}^{2}).\end{split}$ (1.25)
which in turn yields the Hadamard-type formula for the eigenvalue curve $t\mapsto\lambda(t)$ bifurcating from a simple eigenvalue $\lambda_{0}\in\operatorname{Spec}(\mathcal{N}_{t_{0}}+V_{t_{0}})\cap\mathbb{R}$
$\lambda^{\prime}(t_{0})=\frac{\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right)}{\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}},$ (1.26)
where $\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}\neq 0$ follows from $\lambda_{0}\in\mathbb{R}$ and the structure of the canonically symplectic operators.
Our third main result establishes the previous formula via a different method using Lyaponuv-Schmidt reduction and gives an expression for $t^{\prime}(\lambda)$ . In particular, if $\lambda_{0}\in\mathbb{R}$ is an isolated eigenvalue of $\mathcal{N}_{t_{0}}+V_{t_{0}}$ of geometric multiplicity $g$, we show in Proposition 4.6 that the set of points $(\lambda,t)$ near $(\lambda_{0},t_{0})$ such that $\ker(\mathcal{N}_{t}+V_{t}-\lambda)\neq\{0\}$ is given by the zero set of a $g\times g$ matrix $M(\lambda,t)$. We then compute Maslov crossing forms in Lemma 4.5 for the path of Lagrangian planes $\Upsilon_{\lambda,t}\coloneqq\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}$ on the intersection $\left(\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}\right)\cap\mathfrak{D}$, where $\mathbb{K}_{\lambda,t}\coloneqq\operatorname{T}_{t}(\ker((N^{*}+V_{t}^{*}-\lambda)\tau))$, $\mathcal{F}_{t}\coloneqq\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}$ and $\mathfrak{D}$ is the diagonal subspace in $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ (cf. (4.4)),
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})$ $=\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right),$ (1.27a)
$\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})$ $=-\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}.$ (1.27b)
Moreover, we show that in the geometrically simple case (when $g=1$), $t$ and $\lambda$ derivatives of $M$ are given precisely by
$\lambda^{\prime}(t_{0})=-\frac{\mathfrak{m}_{t_{0}}}{\mathfrak{m}_{\lambda_{0}}}\qquad\text{and}\qquad t^{\prime}(\lambda_{0})=-\frac{\mathfrak{m}_{\lambda_{0}}}{\mathfrak{m}_{t_{0}}}\,\,\,\text{if}\,\,\,\mathfrak{m}_{t_{0}}\neq 0.$ (1.28)
We stress that first equation above together with (1.27a), (1.27b) reaffirm the Hadamard-type formula (1.26).
The paper is organised as follows. In Section 2, we describe our abstract set-up and main assumptions, give the Green identity which underpins our analysis, and prove the new symplectic resolvent difference formula for canonically symplectic operators. In Section 3 we use our resolvent difference formula to give a first order expansion of the resolvent, and use this to prove an Hadamard formula for eigenvalue curves $\lambda(t)$. In Section 4 we compute expressions for the Maslov crossing forms, and use Lyaponuv–Schmidt reduction and the implicit function theorem to compute Hadamard formulas for the eigenvalue curves (namely for both $\lambda(t)$ and $t(\lambda)$). In Section 5 we use our expressions for the crossing forms to prove Theorem 1.2 for the example outlined in the introduction.
2. Set-up, symplectic resolvent difference formula and modified Green’s identity
Let $\mathcal{H},\mathfrak{H}$ be separable Hilbert spaces, and let $A$ be a closed, densely defined, symmetric linear operator acting in $\mathcal{H}$ with equal and finite defect indices,
$\begin{split}\dim\ker(A^{*}+i)=\dim\ker(A^{*}-i)<\infty.\end{split}$ (2.1)
We denote the domain of the adjoint $A^{*}$ by $\mathcal{H}_{+}\coloneqq\operatorname{dom}(A^{*})\subset\mathcal{H}$, which is a Hilbert space when equipped with the graph scalar product of $A^{*}$,
$\langle u,v\rangle_{\mathcal{H}_{+}}\coloneqq\langle u,v\rangle_{\mathcal{H}}+\langle A^{*}u,A^{*}v\rangle_{\mathcal{H}}.$ (2.2)
Throughout the rest of the paper we work under the following assumption.
Hypothesis 2.1.
We assume that $A$ is a densely defined, closed, symmetric linear operator acting in $\mathcal{H}$ with equal and finite deficiency indices. Moreover, let
$\operatorname{tr}\coloneqq(\Gamma_{0},\Gamma_{1})^{\top}:\mathcal{H}_{+}\longrightarrow\mathfrak{H}\times\mathfrak{H},$
be a bounded and surjective linear operator, called the trace operator, which satisfies the following abstract Green’s identity for all $u,v\in\operatorname{dom}(A^{*})$,
$\langle A^{*}u,v\rangle_{\mathcal{H}}-\langle u,A^{*}v\rangle_{\mathcal{H}}=\langle\Gamma_{1}u,\Gamma_{0}v\rangle_{\mathfrak{H}}-\langle\Gamma_{0}u,\Gamma_{1}v\rangle_{\mathfrak{H}}.$ (2.3)
Thus, under Hypothesis 2.1, the triple $(\mathfrak{H},\Gamma_{0},\Gamma_{1})$ is a boundary triplet for the adjoint operator $A^{*}$. Consequently, by [Behrndt_Hassi_deSnoo, Proposition 14.5], we have
$\dim\mathfrak{H}=\dim\ker(A^{*}\pm i).$ (2.4)
The right hand side of the identity (2.3) gives rise to a symplectic form $\omega:\mathfrak{H}^{2}\times\mathfrak{H}^{2}\to\mathbb{C}$ defined by
[…]
Thus (2.3) may be written as
$\langle A^{*}u,v\rangle_{\mathcal{H}}-\langle u,A^{*}v\rangle_{\mathcal{H}}=\omega(\operatorname{tr}u,\operatorname{tr}v).$ (2.6)
We denote the annihilator of a subspace $\mathcal{F}\subset\mathfrak{H}\times\mathfrak{H}$ by
$\mathcal{F}^{\circ}\coloneqq\{(f_{1},f_{2})^{\top}\in\mathfrak{H}\times\mathfrak{H}:\omega\left((f_{1},f_{2})^{\top},(g_{1},g_{2})^{\top}\right)=0\,\,\text{for all}\,\,(g_{1},g_{2})^{\top}\in\mathfrak{H}\times\mathfrak{H}\},$ (2.7)
and recall that $\mathcal{F}$ is called Lagrangian if $\mathcal{F}=\mathcal{F}^{\circ}$. We denote by $\Lambda(\mathfrak{H}\times\mathfrak{H})$ the metric space of all Lagrangian subspaces of $\mathfrak{H}\times\mathfrak{H}$ equipped with the metric
$d(\mathcal{F}_{1},\mathcal{F}_{2})\coloneqq\|Q_{1}-Q_{2}\|_{\mathcal{B}(\mathfrak{H}\times\mathfrak{H})},\quad\mathcal{F}_{1},\mathcal{F}_{2}\in\Lambda(\mathfrak{H}\times\mathfrak{H}),$ (2.8)
where $Q_{j}$ is the orthogonal projection onto $\mathcal{F}_{j}$ acting in $\mathfrak{H}\times\mathfrak{H},j=1,2$. Since the deficiency indices of $A$ are finite, there is a one-to-one correspondence between self-adjoint extensions of $A$ the the Lagrangian planes in $\Lambda(\mathfrak{H}\times\mathfrak{H})$, which will be of great importance in the sequel.
We are ready to introduce the operator of interest. Our study takes place on the Hilbert spaces
[…]
Associated with the minimal symmetric operator $A$ is the following closed, densely-defined minimal operator $N$ acting in $\mathcal{H}^{2}$,
$N:\operatorname{dom}(N)=\operatorname{dom}(A)\times\operatorname{dom}(A)\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad N=\begin{pmatrix}0&-A\\ A&0\end{pmatrix}$. (2.9)
The associated adjoint operator $N^{*}$ is given by
$N^{*}:\operatorname{dom}(N^{*})=\mathcal{H}_{+}^{2}\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad N^{*}=\begin{pmatrix}0&A^{*}\\ -A^{*}&0\end{pmatrix}$. (2.10)
As an aside, we note that $\mathcal{H}_{+}^{2}$ is a Hilbert space which we equip the graph scalar product of $N^{*}$,
$\begin{split}\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{H}_{+}^{2}}&\coloneqq\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{H}^{2}}+\langle N^{*}\mathbf{u},N^{*}\mathbf{v}\rangle_{\mathcal{H}^{2}},\\ &=\langle u_{1},v_{1}\rangle_{\mathcal{H}}+\langle u_{2},v_{2}\rangle_{\mathcal{H}}+\langle A^{*}u_{2},A^{*}v_{2}\rangle_{\mathcal{H}}+\langle A^{*}u_{1},A^{*}v_{1}\rangle_{\mathcal{H}}\end{split}$ (2.11)
where $\mathbf{u}=(u_{1},u_{2})^{\top},\mathbf{v}=(v_{1},v_{2})^{\top}\in\mathcal{H}_{+}^{2}$.
Let $\mathcal{A},\mathcal{B}$ be self-adjoint extensions of the minimal symmetric operator $A$, i.e. $A\subset\mathcal{A},\mathcal{B}\subset A^{*}$. The extensions of $N$ whose real spectra we wish to study will be denoted
$\mathcal{N}:\operatorname{dom}(\mathcal{N})=\operatorname{dom}(\mathcal{A})\times\operatorname{dom}(\mathcal{B})\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad\mathcal{N}=\begin{pmatrix}0&-\mathcal{B}\\ \mathcal{A}&0\end{pmatrix}$. (2.12)
On the product space $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ we define the following symplectic form,
$\begin{split}\Omega:\mathfrak{H}^{4}&\times\mathfrak{H}^{4}\longrightarrow\mathbb{C},\\ \Omega((f_{1},&f_{2},f_{3},f_{4})^{\top},(g_{1},g_{2},g_{3},g_{4})^{\top})\coloneqq\langle f_{2},g_{1}\rangle_{\mathfrak{H}}-\langle f_{1},g_{2}\rangle_{\mathfrak{H}}-\langle f_{4},g_{3}\rangle_{\mathfrak{H}}+\langle f_{3},g_{4}\rangle_{\mathfrak{H}},\\ &=\left\langle\mathcal{J}(f_{1},f_{2},f_{3},f_{4})^{\top},(g_{1},g_{2},g_{3},g_{4})^{\top}\right\rangle_{\mathfrak{H}^{4}},\quad\mathcal{J}=\begin{pmatrix}J&0\\ 0&-J\end{pmatrix}=J\oplus(-J),\end{split}$ (2.13)
where $J$ is defined in (2.5). An important operator in our analysis will be the bounded involution
$\tau:\mathcal{H}^{2}\to\mathcal{H}^{2},\quad\tau=\begin{pmatrix}0&I_{\mathcal{H}}\\ I_{\mathcal{H}}&0\end{pmatrix},\quad\tau\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}=\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}$. (2.14)
The operator
$\tau N:\operatorname{dom}(N)=\operatorname{dom}(A)\times\operatorname{dom}(A)\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad\tau N=\begin{pmatrix}A&0\\ 0&-A\end{pmatrix}$, (2.15)
is then symmetric since $\tau N=A\oplus(-A)\subset A^{*}\oplus(-A^{*})=(\tau N)^{*}$, while
$\tau\mathcal{N}:\operatorname{dom}(\mathcal{A})\times\operatorname{dom}(\mathcal{B})\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\quad\tau\mathcal{N}=\begin{pmatrix}\mathcal{A}&0\\ 0&-\mathcal{B}\end{pmatrix}$, (2.16)
is obviously self-adjoint. We note that taking the adjoint is equivalent to conjugating by $\tau$,
$\tau\mathcal{N}^{*}\tau=\mathcal{N}\iff\tau\mathcal{N}\tau=\mathcal{N}^{*}.$ (2.17)
Under Hypothesis 2.1, one can define a trace operator on the product space $\mathcal{H}_{+}\times\mathcal{H}_{+}$ as follows.
Proposition 2.2.
Under Hypothesis 2.1, the linear operator
$\begin{split}&\operatorname{T}\coloneqq\operatorname{tr}\oplus\operatorname{tr}=[\Gamma_{0},\Gamma_{1}]^{\top}\oplus[\Gamma_{0},\Gamma_{1}]^{\top}:\mathcal{H}_{+}^{2}\longrightarrow\mathfrak{H}^{4},\\ &\operatorname{T}\mathbf{u}=\left[\operatorname{tr}u_{1},\operatorname{tr}u_{2}\right]^{\top}=\left[\Gamma_{0}u_{1},\Gamma_{1}u_{1},\Gamma_{0}u_{2},\Gamma_{1}u_{2}\right]^{\top},\quad\mathbf{u}=(u_{1},u_{2})\in\mathcal{H}_{+}^{2},\end{split}$ (2.18)
acting from $\mathcal{H}_{+}^{2}$ into $\mathfrak{H}^{4}$ is bounded and surjective, and satisfies
$\ker(\operatorname{T})=\operatorname{dom}(N)=\operatorname{dom}(A)\times\operatorname{dom}(A).$ (2.19)
It is a trace operator, in the sense that it satisfies the following abstract Green’s identity associated with the maximal adjoint operator $N^{*}$ for functions $\mathbf{u},\mathbf{v}\in\operatorname{dom}\left((\tau N)^{*}\right)=\mathcal{H}_{+}^{2}$,
$\left\langle(\tau N)^{*}\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle\mathbf{u},(\tau N)^{*}\mathbf{v}\right\rangle_{\mathcal{H}^{2}}=\Omega(\operatorname{T}\mathbf{u},\operatorname{T}\mathbf{v}).$ (2.20)
Moreover, for $\lambda\in\mathbb{R}$ we have
$\left\langle(N^{*}-\lambda)\tau\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle\mathbf{u},(N^{*}-\lambda)\tau\mathbf{v}\right\rangle_{\mathcal{H}}=\Omega(\operatorname{T}\mathbf{u},\operatorname{T}\mathbf{v}).$ (2.21)
Proof.
The Green’s identity (2.20) follows from (2.6). Consequently (2.19) follows from the fact that $\operatorname{T}$ is the trace map for symmetric operator $\tau N$ in the Hilbert space $\mathcal{H}\oplus\mathcal{H}$. The identity (2.21) follows from (2.20) and $\tau^{*}=\tau$, $\lambda\in\mathbb{R}$. ∎
In the classical setting of self-adjoint extensions $\mathcal{A}$ of a symmetric operator $A$, the Green identities (2.3) for the adjoint operator $A^{*}$ facilitates the study of the eigenvalues of the self-adjoint restrictions of $A^{*}$ in terms of Lagrangian planes. In the present setting, one immediate impediment to such an approach with the canonical symplectic operator $N$ is that $N$ is not symmetric, i.e. a restriction of its adjoint. Nonetheless, “untwisting” the operator via multiplication by $\tau$ reveals an operator that is symmetric, and for that operator one has the modified Green’s identity (2.20). The restriction that $\lambda\in\mathbb{R}$ and the fact that $\tau$ is symmetric allows one to add in the extra terms involving $\lambda$ in (2.21), and it is this identity that furnishes a symplectic interpretation of the eigenvalues of the extensions (2.12) of the minimal operator $N$.
Next, we write down a symplectic resolvent difference formula for the operator $\mathcal{N}$.
Theorem 2.3.
Assume Hypothesis 2.1, and let $\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{B}_{1},\mathcal{B}_{2}$ be self-adjoint extensions of $A$ with
$\operatorname{tr}(\operatorname{dom}\mathcal{A}_{i})=\operatorname{ran}P_{i},\qquad\operatorname{tr}(\operatorname{dom}\mathcal{B}_{i})=\operatorname{ran}Q_{i},\qquad i=1,2.$ (2.22)
Define $\mathcal{N}_{i}$ by
$\mathcal{N}_{i}:\operatorname{dom}(\mathcal{A}_{i})\times\operatorname{dom}(\mathcal{B}_{i})\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad\mathcal{N}_{i}=\begin{pmatrix}0&-\mathcal{B}_{i}\\ \mathcal{A}_{i}&0\end{pmatrix},\qquad i=1,2.$ (2.23)
Let $\zeta\notin\operatorname{Spec}(\mathcal{N}_{1})\cup\operatorname{Spec}(\mathcal{N}_{2})$ and denote $R_{i}(\zeta)\coloneqq(\mathcal{N}_{i}-\zeta)^{-1},i=1,2$. Then $\bar{\zeta}\notin\operatorname{Spec}(\mathcal{N}_{1})\cup\operatorname{Spec}(\mathcal{N}_{2})$ and one has
$R_{1}(\zeta)-R_{2}(\zeta)=\tau\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}R_{2}(\zeta),$ (2.24)
and
$R_{1}(\zeta)-R_{2}(\zeta)=\tau\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\Big((P_{1}JP_{2})\oplus\left(-Q_{1}JQ_{2}\right)\Big)\operatorname{T}R_{2}(\zeta).$ (2.25)
Proof.
First, we note that $\bar{\zeta}\notin\operatorname{Spec}(\mathcal{N}_{1})\cup\operatorname{Spec}(\mathcal{N}_{2})$ follows from the four fold symmetry of the spectrum of $\mathcal{N}_{i}$. Letting $\mathbf{u},\mathbf{v}\in\mathcal{H}^{2}$ be arbitrary, we have, using that $\tau\mathcal{N}_{1}$ is self-adjoint,
$\left\langle\tau R_{1}(\zeta)\mathbf{u}-\tau R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}=\left\langle R_{1}(\zeta)\mathbf{u}-R_{2}(\zeta)\mathbf{u},\tau(\mathcal{N}_{1}-\bar{\zeta})R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau(\mathcal{N}_{t}-{\zeta})R_{1}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle R_{2}(\zeta)\mathbf{u},\tau\mathcal{N}_{t}R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\zeta\tau R_{2}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle R_{2}(\zeta)\mathbf{u},(\tau N)^{*}R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\zeta\tau R_{2}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle(\tau N)^{*}R_{2}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\zeta\tau R_{2}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+$
$\qquad\qquad\qquad\qquad\Omega(\operatorname{T}R_{2}(\zeta)\mathbf{u},\operatorname{T}R_{1}(\bar{\zeta})\mathbf{v}),$
$\qquad=\left\langle\tau\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle\tau\left(\mathcal{N}_{2}-{\zeta}\right)R_{2}(\zeta)\mathbf{u},R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\mathcal{J}\operatorname{T}R_{2}(\zeta)\mathbf{u},\operatorname{T}R_{1}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathfrak{H}^{4}},$
$\qquad=\left\langle\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
where we used the modified Green’s identity (2.20) in the fourth line, and that
$\tau\mathcal{N}_{1}R_{1}({\zeta})=N^{*}\tau R_{1}({\zeta})$ (2.26)
because $\tau\mathcal{N}_{1}\subset(\tau N)^{*}$ and $\operatorname{ran}(R_{1}({\zeta}))\subset\operatorname{dom}\tau\mathcal{N}_{1}$. This yields (2.24). Next, since $R_{1}(\zeta)\mathbf{u}\in\operatorname{dom}(\mathcal{A}_{1})\times\operatorname{dom}(\mathcal{B}_{1})$, we have
$\operatorname{T}R_{1}(\zeta)\mathbf{u}=\left(P_{1}\oplus Q_{1}\right)\operatorname{T}R_{1}(\zeta)\mathbf{u},$ (2.27)
where we have denoted
$P_{1}\oplus Q_{1}=\begin{bmatrix}P_{1}&0\\ 0&Q_{1}\end{bmatrix}$. (2.28)
Thus
$\left\langle\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}=\left\langle\left(\left(P_{1}\oplus Q_{1}\right)\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\mathcal{J}\left(P_{2}\oplus Q_{2}\right)\operatorname{T}R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad\qquad\qquad=\left\langle\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\left(P_{1}\oplus Q_{1}\right)\mathcal{J}\left(P_{2}\oplus Q_{2}\right)\operatorname{T}R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad\qquad\qquad=\left\langle\left(\operatorname{T}R_{1}(\bar{\zeta})\right)^{*}\Big(\left(P_{1}JP_{2}\right)\oplus\left(-Q_{1}JQ_{2}\right)\Big)\operatorname{T}R_{2}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
from which (2.25) follows. ∎
3. First order asymptotic perturbation theory
In this section, we introduce $t$-dependence of the various operators, working under the following assumptions.
Hypothesis 3.1.
Let
$\operatorname{tr}:[0,1]\to\mathcal{B}(\mathcal{H}_{+},\mathfrak{H}\times\mathfrak{H}):t\mapsto\operatorname{tr}_{t}\coloneqq[\Gamma_{0,t},\Gamma_{1,t}]^{\top}$
be a one-parameter family of trace operators, such that $(\mathfrak{H},\Gamma_{0,t},\Gamma_{1,t})$ is a boundary triplet for each $t\in[0,1]$ satisfying Hypothesis 2.1. Let $P:[0,1]\to\mathcal{B}(\mathfrak{H}^{2}),t\mapsto P_{t}$ and $Q:[0,1]\to\mathcal{B}(\mathfrak{H}^{2}),t\mapsto Q_{t}$ be one-parameter families of orthogonal projections, such that $\operatorname{ran}Q_{t},\operatorname{ran}P_{t}\in\Lambda(\mathfrak{H}^{2})$ are Lagrangian planes for each $t\in[0,1]$. Let $A\subset\mathcal{A}_{t},\mathcal{B}_{t}\subset A^{*}$ be families of self-adjoint extensions of $A$ satisfying
$\operatorname{tr}_{t}(\operatorname{dom}\mathcal{A}_{t})=\operatorname{ran}P_{t},\qquad\operatorname{tr}_{t}(\operatorname{dom}\mathcal{B}_{t})=\operatorname{ran}Q_{t}.$ (3.1)
Define the minimal canonical symplectic operator $N$ as in (2.9), and the following one-parameter family of extensions of $N$,
$\mathcal{N}_{t}:\operatorname{dom}(\mathcal{N}_{t})=\operatorname{dom}(\mathcal{A}_{t})\times\operatorname{dom}(\mathcal{B}_{t})\subseteq\mathcal{H}^{2}\to\mathcal{H}^{2},\qquad\mathcal{N}_{t}=\begin{pmatrix}0&-\mathcal{B}_{t}\\ \mathcal{A}_{t}&0\end{pmatrix}$. (3.2)
For $\operatorname{T}\coloneqq\operatorname{tr}\oplus\operatorname{tr}:[0,1]\to\mathcal{B}(\mathcal{H}_{+}^{2},\mathfrak{H}^{4}):t\mapsto\operatorname{T}_{t}=\operatorname{tr}_{t}\oplus\operatorname{tr}_{t}=[\Gamma_{0,t},\Gamma_{1,t}]^{\top}\oplus[\Gamma_{0,t},\Gamma_{1,t}]^{\top},$ we note note
$\operatorname{T}_{t}(\operatorname{dom}\mathcal{N}_{t})=\operatorname{T}_{t}(\operatorname{dom}\mathcal{A}_{t}\times\operatorname{dom}\mathcal{B}_{t})=\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}.$ (3.3)
Next, let $F:[0,1]\to\mathcal{B}(\mathcal{H}),t\mapsto F_{t}$ and $G:[0,1]\to\mathcal{B}(\mathcal{H}),t\mapsto G_{t}$ be families of bounded, self-adjoint operators acting in $\mathcal{H}$. On the product space $\mathcal{H}^{2}$, the mapping
$V:[0,1]\to\mathcal{B}(\mathcal{H}^{2}),\qquad V_{t}=\begin{pmatrix}0&-G_{t}\\ F_{t}&0\end{pmatrix}$, (3.4)
is then a one-parameter family of bounded operators, which, similar to (2.17) satisfies
$V_{t}^{*}=\tau V_{t}\tau\iff V_{t}=\tau V_{t}^{*}\tau.$ (3.5)
For each $t\in[0,1]$, we define the resolvent $R_{t}(\zeta)\coloneqq(\mathcal{N}_{t}+V_{t}-\zeta)^{-1}\in\mathcal{B}(\mathcal{H}^{2})$ for all $\zeta\notin\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$.
Our first task is to write down a resolvent difference formula for the resolvent
$R_{t}(\zeta)=(\mathcal{N}_{t}+V_{t}-\zeta)^{-1},\qquad\zeta\notin\operatorname{Spec}(\mathcal{N}_{t}+V_{t}),$ (3.6)
now incorporating the bounded potential $V_{t}$.
Lemma 3.2.
Let $t,s,r\in[0,1]$ and $\zeta\notin\operatorname{Spec}(\mathcal{N}_{t}+V_{t})\cup\operatorname{Spec}(\mathcal{N}_{s}+V_{s})$. Then for $R_{t}(\zeta)\coloneqq(\mathcal{N}_{t}+V_{t}-\zeta)^{-1}$ we have
$R_{t}(\zeta)$ $-R_{s}(\zeta)=R_{t}({\zeta})(V_{s}-V_{t})R_{s}(\zeta)+\tau\left(\operatorname{T}_{r}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}_{r}R_{s}(\zeta),$ (3.7)
$\begin{split}&=R_{t}({\zeta})(V_{s}-V_{t})R_{s}(\zeta)+\tau\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\Big((P_{t}-P_{s})J\oplus\left(-(Q_{t}-Q_{s})J\right)\Big)\operatorname{T}_{s}R_{s}(\zeta)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{s})R_{s}(\zeta).\end{split}$ (3.8)
Proof.
Let $\mathbf{u},\mathbf{v}\in\mathcal{H}^{2}$ be arbitrary. We have, using that $\tau(\mathcal{N}_{t}+V_{t})$ is self-adjoint,
$\left\langle\tau R_{t}(\zeta)\mathbf{u}-\tau R_{s}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}=\left\langle R_{t}(\zeta)\mathbf{u}-R_{s}(\zeta)\mathbf{u},\tau(\mathcal{N}_{t}+V_{t}-\bar{\zeta})R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau(\mathcal{N}_{t}+V_{t}-{\zeta})R_{t}(\zeta)\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle R_{s}(\zeta)\mathbf{u},\tau(\mathcal{N}_{t}+V_{t}-\bar{\zeta})R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle R_{s}(\zeta)\mathbf{u},\tau(V_{s}-V_{t})R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}-\left\langle R_{s}(\zeta)\mathbf{u},((\tau N+\tau V_{s})^{*}-\tau\bar{\zeta})R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad=\left\langle\tau\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle R_{s}(\zeta)\mathbf{u},\tau(V_{s}-V_{t})R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}$
$\qquad\qquad-\left\langle((\tau N+\tau V_{s})^{*}-\tau{\zeta})R_{s}(\zeta)\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\Omega(\operatorname{T}_{r}R_{s}(\zeta)\mathbf{u},\operatorname{T}_{r}R_{t}(\bar{\zeta})\mathbf{v}),$
$\qquad=\left\langle\tau\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle R_{s}(\zeta)\mathbf{u},\tau(V_{s}-V_{t})\tau^{2}R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
$\qquad\qquad-\left\langle\tau(\mathcal{N}_{s}+V_{s}-{\zeta})R_{s}(\zeta)\mathbf{u},R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\mathcal{J}\operatorname{T}_{r}R_{s}(\zeta)\mathbf{u},\operatorname{T}_{r}R_{t}(\bar{\zeta})\mathbf{v}\right\rangle_{\mathfrak{H}^{4}},$
$\qquad=\left\langle\tau R_{t}(\zeta)\tau(V_{s}-V_{t})^{*}\tau R_{s}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}}+\left\langle\left(\operatorname{T}_{r}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}_{r}R_{s}(\zeta)\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{H}^{2}},$
which, using (3.5), proves (3.7). (Note that we used (2.20) with trace operator $\operatorname{T}_{r}$.)
Next, using (2.27) and that $P_{s}JP_{s}=Q_{s}JQ_{s}=0$ since $\operatorname{ran}P_{s}$ and $\operatorname{ran}Q_{s}$ are Lagrangian, we have,
$\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}_{t}R_{s}(\zeta)=\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}\operatorname{T}_{s}R_{s}(\zeta)+\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{s})R_{s}(\zeta),$
$\qquad=\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}(P_{t}\oplus Q_{t})\mathcal{J}(P_{s}\oplus Q_{s})\operatorname{T}_{s}R_{s}(\zeta)+\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{s})R_{s}(\zeta),$
$\qquad=\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}(P_{t}JP_{s}\oplus(-Q_{t}JQ_{s})\operatorname{T}_{s}R_{s}(\zeta)+\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{s})R_{s}(\zeta),$
$\qquad=\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\left((P_{t}-P_{s})J\oplus(-(Q_{t}-Q_{s})J)\right)\operatorname{T}_{s}R_{s}(\zeta)+\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{s})R_{s}(\zeta),$
which yields (LABEL:resolvent_diff2new) upon substituting $r=t$ into (3.7). ∎
3.1. Asymptotic expansions and a Hadamard formula
In this section, we give first order asymptotic expansions for the resolvent, and for an operator to which $P(t)(\mathcal{N}_{t}+V_{t})P(t)$, where $P(t)$ is the spectral projection onto the finite dimensional kernel $\mathcal{N}_{t}+V_{t}-\lambda$, is similar. Using the latter, we give a Hadamard-type variational formula for the eigenvalue curve $\lambda(t)$.
Hypothesis 3.3.
Let us fix $t_{0}\in(0,1)$ and suppose that there exists $z\in\mathbb{C}$ such that $z\not\in\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$ for $t$ near $t_{0}$.
Next, we give two auxiliary statements regarding boundedness and continuity of the mapping $t\mapsto(\mathcal{N}_{t}+V_{t}-z)^{-1}$. The first of the following statements will be required when applying the trace operator $\operatorname{T}_{t}$ to the resolvent $R_{t}(\zeta)$, where $\operatorname{dom}(\operatorname{T}_{t})=\mathcal{H}_{+}^{2}$ (recall Proposition 2.2). In the proof of Theorem 3.5, we will also need to control the operator norm of the resolvent, viewing it as a bounded operator from $\mathcal{H}^{2}$ into $\mathcal{H}_{+}^{2}$. Second, we prove the continuity of the resolvent operator in $t$, viewed as a bounded operator from $\mathcal{H}^{2}$ into $\mathcal{H}_{+}^{2}$.
Proposition 3.4.
Assume Hypothesis 3.1 and Hypothesis 3.3. Then the following assertions hold.
  1. (1)
    For $t$ near $t_{0}$, the resolvent $(\mathcal{N}_{t}+V_{t}-z)^{-1}$ can be viewed as a bounded operator from $\mathcal{H}^{2}$ into $\mathcal{H}_{+}^{2}$. Consequently, $\operatorname{T}_{t}(\mathcal{N}_{t}+V_{t}-z)^{-1}\in\mathcal{B}(\mathcal{H}^{2},\mathfrak{H}^{4})$.
  2. (2)
    If the mappings $t\mapsto P_{t},t\mapsto Q_{t},t\mapsto\operatorname{T}_{t}$ are continuous at $t_{0}$, then so is the mapping $t\mapsto(\mathcal{N}_{t}+V_{t}-z)^{-1}\in\mathcal{B}(\mathcal{H}^{2},\mathcal{H}^{2}_{+})$, that is,
    $\|(\mathcal{N}_{t}+V_{t}-z)^{-1}-(\mathcal{N}_{t_{0}}+V_{t_{0}}-z)^{-1}\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}=o(1),\,\,\,t\to t_{0}.$ (3.9)
Proof.
For (1), recall that $\mathcal{H}_{+}^{2}$ is equipped with the norm (2.11). Then for all $\mathbf{u}\in\mathcal{H}^{2}$, using the Cauchy-Schwartz inequality and that $\tau\mathcal{N}_{t}\subset(\tau N)^{*}$, we have
$\|R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}_{+}^{2}}=\|\tau R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}_{+}^{2}}=\|\tau R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}}+\|N^{*}\tau R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}},$
$\quad=\|\tau R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}}+\|(\tau N)^{*}R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}},$
$\quad\leq\|\tau R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}}+\left(\|\tau(\mathcal{N}_{t}+V_{t}-z)R_{t}(z)\mathbf{u}\|_{\mathcal{H}^{2}}+(\|\tau V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}+|z|)\|\tau R_{t}(z)\mathbf{u}\|_{\mathcal{H}^{2}}\right)^{2},$
$\quad=\|R_{t}(z)\mathbf{u}\|^{2}_{\mathcal{H}^{2}}+\left(\|\mathbf{u}\|_{\mathcal{H}^{2}}+(\|\tau V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}+|z|)\|R_{t}(z)\mathbf{u}\|_{\mathcal{H}^{2}}\right)^{2},$
it follows that
$\|R_{t}(z)\|^{2}_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}\leq\|R_{t}(z)\|^{2}_{\mathcal{B}(\mathcal{H}^{2})}+\left(1+(\|\tau V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}+|z|)\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2})}\right)^{2},$ (3.10)
and, hence,
$\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}\leq 1+(\|\tau V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}+|z|+1)\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2})}.$ (3.11)
Regarding statement (2), we first claim that
$\left\|R_{t}(z)-R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}\leq\sqrt{1+|z|}\left\|R_{t}(z)-R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}.$ (3.12)
Indeed, again using that $\tau\mathcal{N}_{t}\subset(\tau N)^{*}$, for $\mathbf{u}\in\mathcal{H}^{2}$ we have
$\left\|R_{t}(z)\mathbf{u}-R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}_{+}^{2}}=\left\|\tau R_{t}(z)\mathbf{u}-\tau R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}_{+}^{2}},$
$\qquad=\left\|\tau R_{t}(z)\mathbf{u}-\tau R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}}+\left\|(\tau N)^{*}R_{t}(z)\mathbf{u}-(\tau N)^{*}R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}},$
$\qquad=\left\|\tau R_{t}(z)\mathbf{u}-\tau R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}}$
$\qquad\quad+\left\|\tau(\mathcal{N}_{t}+V_{t}-z)R_{t}(z)\mathbf{u}-(\tau(\mathcal{N}_{t_{0}}+V_{t_{0}}-z)R_{t_{0}}(z)\mathbf{u}-\tau(V_{t}-z)R_{t}(z)\mathbf{u}+\tau(V_{t_{0}}-z)R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}},$
$\qquad=\left\|R_{t}(z)\mathbf{u}-R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}}+\left\|\tau(V_{t}-z)R_{t}(z)\mathbf{u}-\tau(V_{t_{0}}-z)R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}},$
$\qquad\leq\left\|R_{t}(z)\mathbf{u}-R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}}+2\left\|\tau(V_{t}-z)\right\|^{2}_{\mathcal{B}(\mathcal{H}^{2})}\left\|(R_{t}(z)-R_{t_{0}}(z))\mathbf{u}\|^{2}_{\mathcal{H}^{2}}+2\|\tau(V_{t}-V_{t_{0}})\|^{2}_{\mathcal{B}(\mathcal{H}^{2})}\|R_{t_{0}}(z)\mathbf{u}\right\|^{2}_{\mathcal{H}^{2}},$
$\qquad\leq\left(\|R_{t}(z)\mathbf{u}-R_{t_{0}}(z)\mathbf{u}\|_{\mathcal{H}^{2}}\left(1+\sqrt{2}\left\|\tau(V_{t}-z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}\right)+\sqrt{2}\|\tau(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}\left\|R_{t_{0}}(z)\mathbf{u}\right\|_{\mathcal{H}^{2}}\right)^{2}$
Taking the supremum over all $\mathbf{u}\in\mathcal{H}^{2}$ such that $\|\mathbf{u}\|_{\mathcal{H}^{2}}=1$ we obtain
$\left\|R_{t}(z)-R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}$
$\quad\leq\|R_{t}(z)-R_{t_{0}}(z)\|_{\mathcal{B}(\mathcal{H}^{2})}\left(1+\sqrt{2}\left\|\tau(V_{t}-z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}\right)+\sqrt{2}\|\tau(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}\left\|R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}$
Thus, it is enough to prove that the right hand side of (3.12) is $o(1)$ as $t\to t_{0}$. To this end, we first establish an auxiliary assertions
$\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}=\mathcal{O}(1),\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2})}=\mathcal{O}(1),\ t\rightarrow t_{0}.$ (3.13)
Using the resolvent difference formula (LABEL:resolvent_diff2new), we find
$\begin{split}&\left\|R_{t}(z)-R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}\\ &\qquad\leq\|R_{t}({z})(V_{t}-V_{t_{0}})R_{t_{0}}(z)\|_{\mathcal{B}(\mathcal{H}^{2})}+\|\tau\left(\operatorname{T}_{t}R_{t}(\bar{z})\right)^{*}\Big((P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\Big)\mathcal{J}T_{t_{0}}R_{t_{0}}(z)\|_{\mathcal{B}(\mathcal{H}^{2})},\\ &\qquad\leq\ \|R_{t}({z})\|_{\mathcal{B}(\mathcal{H}^{2})}\|(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}\|R_{t_{0}}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2})}\\ &\qquad\qquad+\|\operatorname{T}_{t}\|_{\mathcal{B}(\mathcal{H}_{+}^{2},\mathfrak{H}^{4})}\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\|_{\mathcal{B}(\mathfrak{H}^{4})}\times\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left\|\operatorname{T}_{t_{0}}\right\|_{\mathcal{B}(\mathcal{H}_{+}^{2},\mathfrak{H}^{4})}\left\|R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})},\\ &\qquad\leq C(\|R_{t}({z})\|_{\mathcal{B}(\mathcal{H}^{2})}\|(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}+\left\|R_{t}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}\left\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\right\|_{\mathcal{B}(\mathfrak{H}^{4})}),\end{split}$ (3.14)
for some $C=C(z)>0$. Further employing (3.11) and the triangle inequality we arrive at
$\left\|R_{t}(z)-R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}$ (3.15)
$\qquad\leq C(\|R_{t}({\zeta})\|_{\mathcal{B}(\mathcal{H}^{2})}\|(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}+(1+\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2})})\left\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\right\|_{\mathcal{B}(\mathfrak{H}^{4})}),$ (3.16)
for some $C=C(z)>0$. Next, we use the triangle inequality in the right-hand side above and rearrange terms to get
$\left\|R_{t}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}\left(1-C(\|(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}+\left\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\right\|_{\mathcal{B}(\mathfrak{H}^{4})})\right)$ (3.17)
$\qquad\leq\left\|R_{t_{0}}(z)\right\|_{\mathcal{B}(\mathcal{H}^{2})}+C\left\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\right\|_{\mathcal{B}(\mathfrak{H}^{4})},$ (3.18)
which together with
$\|(V_{t}-V_{t_{0}})\|_{\mathcal{B}(\mathcal{H}^{2})}+\left\|(P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}})\right\|_{\mathcal{B}(\mathfrak{H}^{4})}=o(1),\ t\rightarrow t_{0},$ (3.19)
yields the second equality in (3.13). The first equality in (3.13) follows form the second one and (3.11).
Finally, (3.9) follows from (3.12), (3.13), (LABEL:eq:res_dif_norm_bound), (3.19). ∎
In the following theorem we show that the continuity, Lipschitz continuity and differentiability of the mappings $t\mapsto V_{t}$, $t\mapsto\operatorname{T}_{t},t\mapsto Q_{t},t\mapsto P_{t}$ imply the continuity, Lipschitz continuity and differentiability, respectively, of the mapping $t\mapsto R_{t}(\zeta)$.
Theorem 3.5.
Assume Hypothesis 3.3. Suppose that $\zeta_{0}\notin\operatorname{Spec}(\mathcal{N}_{t_{0}}+V_{t_{0}})$ and define
$\mathcal{U}_{\varepsilon}\coloneqq\{(t,\zeta)\in(0,1)\times\mathbb{C}:|t-t_{0}|\leq\varepsilon,|\zeta-\zeta_{0}|\leq\varepsilon\}\,\,\,\text{for}\,\,\,\varepsilon>0.$ (3.20)
  1. (1)
    Suppose that the mappings $t\mapsto V_{t}$, $t\mapsto\operatorname{T}_{t},t\mapsto Q_{t},t\mapsto P_{t}$ are continuous at $t_{0}$. Then there exists an $\varepsilon>0$ such that if $(t,\zeta)\in\mathcal{U}_{\varepsilon}$ then $\zeta\notin\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$, and operator valued function $t\mapsto R_{t}(\zeta)=(\mathcal{N}_{t}+V_{t}-\zeta)^{-1}\in\mathcal{B}(\mathcal{H}^{2})$ is continuous at $t_{0}$ uniformly for $|\zeta-\zeta_{0}|<\varepsilon$.
  2. (2)
    Suppose that the mappings $t\mapsto V_{t}$, $t\mapsto\operatorname{T}_{t},t\mapsto Q_{t},t\mapsto P_{t}$ are Lipschitz continuous at $t_{0}$. Then there exists a constant $c>0$ such that for all $(t,\zeta)\in\mathcal{U}_{\varepsilon}$, we have
    $\|R_{t}(\zeta)-R_{t_{0}}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2})}\leq c|t-t_{0}|.$ (3.21)
  3. (3)
    Suppose that the mappings $t\mapsto V_{t}$, $t\mapsto\operatorname{T}_{t},t\mapsto Q_{t},t\mapsto P_{t}$ are differentiable at $t_{0}$. Then for some $\varepsilon>0$ the following asymptotic expansion holds uniformly for $|\zeta-\zeta_{0}|<\varepsilon$,
    $\begin{split}R_{t}(\zeta)&\underset{t\to t_{0}}{=}R_{t_{0}}(\zeta)+\Big(-R_{t_{0}}(\zeta)\dot{V}_{t_{0}}R_{t_{0}}(\zeta)+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)\\ &\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}R_{t_{0}}(\zeta)\Big)(t-t_{0})+o(t-t_{0})\quad in\,\,\mathcal{B}(\mathcal{H}^{2}).\end{split}$ (3.22)
Proof.
Assertion (1) follows immediately from Proposition 3.4 (2).
For statement (2), by(1) we have
$\sup\left\{\|R_{t}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2})}:(t,\zeta)\in\mathcal{U}_{\varepsilon}\right\}<\infty.$ (3.23)
We claim that yet a smaller choice of $\varepsilon>0$ gives
$\sup\left\{\|R_{t}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}:(t,\zeta)\in\mathcal{U}_{\varepsilon}\right\}<\infty.$ (3.24)
Indeed, by the resolvent identity we have
$R_{t}(\zeta)=R_{t}(z)-(z-\zeta)R_{t}(z)R_{t}(\zeta).$ (3.25)
Using this and the fact that
$\|R_{t}(z)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}=\mathcal{O}(1),\quad t\to t_{0},$
(3.23) then implies (3.24). Next, by the resolvent difference formula (LABEL:resolvent_diff2new) and (3.24), we infer
$R_{t}(\zeta)-R_{t_{0}}(\zeta)$ $=R_{t}(\zeta)(V_{t_{0}}-V_{t})R_{t_{0}}(\zeta)$
$\qquad+\tau(\operatorname{T}_{t}R_{t}(\bar{\zeta}))^{*}((P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}}))\mathcal{J}(P_{t_{0}}\oplus Q_{t_{0}})\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)$
$\qquad\quad+\tau(\operatorname{T}_{t}R_{t}(\bar{\zeta}))^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{t_{0}})R_{t_{0}}(\zeta).$
Hence
$\|R_{t}(\zeta)-R_{t_{0}}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2})}$ $\leq c\max\Big\{\|((P_{t}-P_{t_{0}})\oplus(Q_{t}-Q_{t_{0}}))\|_{\mathcal{B}(\mathfrak{H}^{4})},$
$\qquad\qquad\qquad\qquad\qquad\qquad\|\operatorname{T}_{t}-\operatorname{T}_{t_{0}}\|_{\mathcal{B}(\mathcal{H}_{+}^{2},\mathfrak{H}^{4})},\|V_{t}-V_{t_{0}}\|_{\mathcal{B}(\mathcal{H}^{2})}\Big\},$
for some $c>0$ and all $(t,\zeta)\in\mathcal{U}_{\varepsilon}$, where we used that
$\|\operatorname{T}_{t}R_{t}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2},\mathfrak{H}^{4})}\leq\|\operatorname{T}_{t}\|_{\mathcal{B}(\mathcal{H}_{+}^{2},\mathfrak{H}^{4})}\|R_{t}(\bar{\zeta})\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}.$ (3.26)
For the proof of statement (3), we first note that the mapping $t\mapsto R_{t}(\zeta)\in\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})$ is continuous at $t_{0}$, i.e.
$\|R_{t}(\zeta)-R_{t_{0}}(\zeta)\|_{\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})}=o(1),\quad t\to t_{0},$ (3.27)
uniformly for $|\zeta-\zeta_{0}|<\varepsilon$ for $\varepsilon>0$ as in (3.24). Next, by assumption we have
$\begin{split}Q_{t}&\underset{t\to t_{0}}{=}Q_{t_{0}}+\dot{Q}_{t_{0}}(t-t_{0})+o(t-t_{0});\ P_{t}\underset{t\to t_{0}}{=}P_{t_{0}}+\dot{P}_{t_{0}}(t-t_{0})+o(t-t_{0}),\\ V_{t}&\underset{t\to t_{0}}{=}V_{t_{0}}+\dot{V}_{t_{0}}(t-t_{0})+o(t-t_{0});\ T_{t}\underset{t\to t_{0}}{=}\operatorname{T}_{t_{0}}+\,\dot{\operatorname{T}}_{t_{0}}(t-t_{0})+o(t-t_{0}).\end{split}$ (3.28)
Substituting $s=t_{0}$ into the resolvent difference formula (LABEL:resolvent_diff2new), we have, using (3.21) and (3.28),
$R_{t}(\zeta)$ $-R_{t_{0}}(\zeta)\,=R_{t}({\zeta})\left(V_{t_{0}}-V_{t}\right)R_{t_{0}}(\zeta)+\tau\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\Big((P_{t}-P_{t_{0}})J\oplus\left(-(Q_{t}-Q_{t_{0}})J\right)\Big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t}R_{t}(\bar{\zeta})\right)^{*}\mathcal{J}(\operatorname{T}_{t}-\operatorname{T}_{t_{0}})R_{t_{0}}(\zeta),$
$\hskip-4.2679pt\underset{t\to t_{0}}{=}\left(R_{t_{0}}(\zeta)+\mathcal{O}(t-t_{0})\right)\left(-\dot{V}_{t_{0}}(t-t_{0})+o(t-t_{0})\right)R_{t_{0}}(\zeta)$
$\qquad+\tau\left(\left(\operatorname{T}_{t_{0}}+\mathcal{O}(t-t_{0})\right)\left(R_{t_{0}}(\bar{\zeta})+\mathcal{O}_{\|\cdot\|_{\mathcal{B}\left(\mathcal{H}^{2},\mathcal{H}_{+}^{2}\right)}}(1)\right)\right)^{*}\times$
$\qquad\Big((\dot{P}_{t_{0}}(t-t_{0})+o(t-t_{0}))J\oplus-(\dot{Q}_{t_{0}}(t-t_{0})+o(t-t_{0}))J\Big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)$
$\qquad+\tau\left(\left(\operatorname{T}_{t_{0}}+\mathcal{O}(t-t_{0})\right)\left(R_{t_{0}}(\bar{\zeta})+\mathcal{O}_{\|\cdot\|_{\mathcal{B}\left(\mathcal{H}^{2},\mathcal{H}_{+}^{2}\right)}}(1)\right)\right)^{*}\mathcal{J}(\dot{\operatorname{T}}_{t_{0}}(t-t_{0})+o(t-t_{0}))R_{t_{0}}(\zeta),$
$\hskip-5.69054pt\underset{t\to t_{0}}{=}\Big(-R_{t_{0}}(\zeta)\dot{V}_{t_{0}}R_{t_{0}}(\zeta)+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)$
$\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}R_{t_{0}}(\zeta)\Big)(t-t_{0})+o(t-t_{0}),$
in $\mathcal{B}(\mathcal{H}^{2})$ uniformly for $|\zeta-\zeta_{0}|<\varepsilon$. This proves (3.22). ∎
Hypothesis 3.6.
At a given $t_{0}\in[0,1]$, we suppose that $\lambda$ is an isolated, simple eigenvalue of $\mathcal{N}_{t_{0}}+V_{t_{0}}$. We denote by $\gamma$ the following contour surrounding $\lambda$,
[…]
By Theorem 3.5 for $t$ near $t_{0}$ the contour $\gamma$ encloses simple eigenvalue $\lambda(t)\in\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$. Corresponding to this eigenvalue is the Riesz projection
$P(t)\coloneqq-\frac{1}{2\pi i}\int_{\gamma}R_{t}(\zeta)d\zeta,\qquad R_{t}(\zeta)=(\mathcal{N}_{t}+V_{t}-\zeta)^{-1},$ (3.30)
where $\gamma$ has the anti-clockwise orientation, as well as the reduced resolvent
$S\coloneqq\frac{1}{2\pi i}\int_{\gamma}(\zeta-\lambda)^{-1}R_{t_{0}}(\zeta)d\zeta,$ (3.31)
and the identity
$P(t_{0})R_{t_{0}}(\zeta)=(\lambda-\zeta)^{-1}P(t_{0}).$ (3.32)
Remark 3.7.
We note that (3.32) holds due to our assumption that $\lambda$ is a simple eigenvalue. It would still continue to hold for semi-simple eigenvalues (i.e., those with equal algebraic and geometric multiplicities). However in a more general case of non-semi-simple eigenvalues the right-hand side of (3.32) contains a nilpotent part, cf., e.g., [Kato, eq. (5.23) in Chapter I.5] which would enter an Hadamart-type formula for the derivative of eigenvalue curves.
Remark 3.8.
Due to continuity of the mapping $\zeta\mapsto R_{t}(\zeta)\in\mathcal{B}(\mathcal{H}^{2},\mathcal{H}^{2}_{+})$ we can view $P(t),S$ as operators in $\mathcal{B}(\mathcal{H}^{2},\mathcal{H}_{+}^{2})$ and we have
$\begin{split}\frac{1}{2\pi i}\int_{\gamma}\operatorname{T}_{t}\left((\zeta-\lambda)^{-1}R_{t}(\zeta)\right)d\zeta&=\operatorname{T}_{t}\frac{1}{2\pi i}\int_{\gamma}\left((\zeta-\lambda)^{-1}R_{t}(\zeta)\right)d\zeta=\operatorname{T}_{t}S,\\ (\operatorname{T}_{t}P(t))&\in\mathcal{B}(\mathcal{H}^{2},\mathfrak{H}^{4}).\end{split}$ (3.33)
We remark that, since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, $P(t)$ and $S$ are not self-adjoint, see for example [Gohberg_Goldberg_Kaashoek, Proposition I.2.5]. However, using (2.17) and (3.5) we have
$(\tau P(t))^{*}=P(t)^{*}\tau=\frac{1}{2\pi i}\int_{\bar{\gamma}}(\mathcal{N}_{t}^{*}+V_{t}^{*}-\zeta)^{-1}\tau\,d\zeta=\frac{-1}{2\pi i}\int_{\gamma}\tau(\mathcal{N}_{t}+V_{t}-\zeta)^{-1}d\zeta=\tau P(t),$ (3.34)
and, similarly, $(\tau S)^{*}=S^{*}\tau=\tau S$.
Towards a Hadamard formula, we need to write down an asymptotic expansion for $P(t)H_{t}P(t)$ for $t$ near $t_{0}$. Since the domain of $P(t)H_{t}P(t)$ is $t$-dependent, we introduce the following transformation operators which allow us to work with a similar operator for which the domain is fixed, cf. [Kato, Section I.4.6]. To that end, we define $D(t)\coloneqq P(t)-P(t_{0})$ satisfying $\|D(t)\|_{\mathcal{B}(\mathcal{H})}=o(1)$ as $t\to t_{0}$, which follows from (3.21) and the definition of $P(t)$ (3.30). The following operators are then well defined for $t$ near $t_{0}$.
$\begin{split}U(t)&\coloneqq(I-D^{2}(t))^{-1/2}\Big((I-P(t))(I-P(t_{0}))+P(t)P(t_{0})\Big),\\ U(t)^{-1}&=\Big((I-P(t_{0}))(I-P(t))+P(t_{0})P(t)\Big)(I-D^{2}(t))^{-1/2}.\end{split}$ (3.35)
Moreover, as in [Kato, Section I.4.6] and [F04, Proposition 2.18], we have that
$U(t)P(t_{0})=P(t)U(t),$ (3.36)
and that $U(t)$ maps $\operatorname{ran}P(t_{0})$ onto $\operatorname{ran}P(t)$ for $t$ near $t_{0}$. Unlike the analysis in [LS20first], since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, the Riesz projection $P(t)$ is not self-adjoint and therefore not orthogonal, and hence $U(t)$ is not unitary. (In fact, one can show that $U(t)^{-1}=\tau U(t)^{*}\tau$.) Nonetheless, only minor tweaks to the proof of [LS20first, Lemma 3.24] are required for the following lemma.
Lemma 3.9.
For a given $t_{0}$, we assume that the mappings $t\mapsto V_{t}$, $t\mapsto\operatorname{T}_{t},t\mapsto Q_{t},t\mapsto P_{t}$ are differentiable at $t_{0}$, and that Hypothesis 3.6 holds. We then have
$\begin{split}&P(t_{0})U(t)^{-1}(\mathcal{N}_{t}+V_{t})P(t)U(t)P(t_{0})\underset{t\to t_{0}}{=}\lambda P(t_{0})+\Big(P(t_{0})\dot{V}_{t_{0}}P(t_{0})\\ &\,-\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\left(\dot{Q}_{t_{0}}\oplus\dot{P}_{t_{0}}\right)\mathcal{J}\operatorname{T}_{t_{0}}P(t_{0})-\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0})\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.37)
Proof.
We expand the left hand side by making use of the resolvent expansion (3.22) for $t$ near $t_{0}$. Multiplying (3.22) on the right by $P(t_{0})$ and using the identity
$R_{t_{0}}(\zeta)P(t_{0})=P(t_{0})R_{t_{0}}(\zeta)=(\lambda-\zeta)^{-1}P(t_{0}),$ (3.38)
we find that
$\begin{split}R_{t}(\zeta)P(t_{0})\underset{t\to t_{0}}{=}&(\lambda-\zeta)^{-1}P(t_{0})+(\lambda-\zeta)^{-1}\Big(-R_{t_{0}}(\zeta)\dot{V}_{t_{0}}P(t_{0})\\ &\qquad+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}P(t_{0})\\ &\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0})\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.39)
Similarly, multiplying (3.22) on the left by $P(t_{0})$, using (3.38) and that $P(t_{0})\tau=\tau P(t_{0})^{*}$, we have
$\begin{split}P(t_{0})R_{t}(\zeta)\underset{t\to t_{0}}{=}&(\lambda-\zeta)^{-1}P(t_{0})+(\lambda-\zeta)^{-1}\Big(-P(t_{0})\dot{V}_{t_{0}}R_{t_{0}}(\zeta)\\ &\qquad+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)\\ &\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}R_{t_{0}}(\zeta)\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.40)
The proof is split into several steps.
Step 1. We have
$P(t_{0})P(t)P(t_{0})\underset{t\to t_{0}}{=}P(t_{0})+o(t-t_{0}).$ (3.41)
Proof.
For any continuous $F:\gamma\to\mathcal{B}(\mathcal{H}^{2},\mathfrak{H}^{4})$ we have
$\left(\int_{\gamma}F(\zeta)d\zeta\right)^{*}=-\int_{\gamma}(F(\bar{\zeta}))^{*}d\zeta.$
Applying this to $F(\zeta)=\frac{1}{2\pi i}(\lambda-\zeta)^{-1}\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)$, and using (3.31), (3.33), we have
$\int_{\gamma}\left(\frac{1}{2\pi i}(\lambda-\bar{\zeta})^{-1}\operatorname{T}_{t_{0}}R_{t_{0}}(\bar{\zeta})\right)^{*}d\zeta=\left(-\int_{\gamma}\frac{1}{2\pi i}(\lambda-\zeta)^{-1}\operatorname{T}_{t_{0}}R_{t_{0}}(\zeta)d\zeta\right)^{*}=(\operatorname{T}_{t_{0}}S)^{*}.$
Using this, multiplying both sides of (3.39) by $\frac{1}{2\pi i}$ and integrating over $\gamma$, we obtain
$\begin{split}P(t)P(t_{0})&\underset{t\to t_{0}}{=}P(t_{0})+\Big(-S\dot{V}_{t_{0}}P(t_{0})+\tau\left(\operatorname{T}_{t_{0}}S\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}P(t_{0})\\ &\qquad\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}S\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0})\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.42)
Similarly, multiplying both sides of (3.40) by $\frac{1}{2\pi i}$ and integrating over $\gamma$, we obtain
$\begin{split}P(t_{0})P(t)&\underset{t\to t_{0}}{=}P(t_{0})+\Big(-P(t_{0})\dot{V}_{t_{0}}S+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}S\\ &\qquad\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\operatorname{T}_{t_{0}}S\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.43)
Multiplying on the right by $P(t_{0})$ and using $SP(t_{0})=0$, we arrive at (3.41). ∎
Step 2. We have
$P(t_{0})U(t)P(t_{0})$ $\underset{t\to t_{0}}{=}P(t_{0})+o(t-t_{0}),$ (3.44)
$P(t_{0})U(t)^{-1}P(t_{0})$ $\underset{t\to t_{0}}{=}P(t_{0})+o(t-t_{0}),$ (3.45)
$\begin{split}\left(I-P(t_{0})\right)U(t)P(t_{0})&\underset{t\to t_{0}}{=}(I-P(t_{0}))\Big(-S\dot{V}_{t_{0}}P(t_{0})+\tau\left(\operatorname{T}_{t_{0}}S\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}P(t_{0})\\ &\hskip-40.00006pt\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}S\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0})\Big)(t-t_{0})+o(t-t_{0}),\end{split}$ (3.46)
$\begin{split}P(t_{0})U(t)^{-1}(I-P(t_{0}))&\underset{t\to t_{0}}{=}\Big(-P(t_{0})\dot{V}_{t_{0}}S+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\big(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\big)\operatorname{T}_{t_{0}}S\\ &\hskip-40.00006pt\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}S\Big)(I-P(t_{0}))(t-t_{0})+o(t-t_{0}).\end{split}$ (3.47)
Proof.
Using (3.30) and (3.21), we have
$D(t)=P(t)-P(t_{0})=-\frac{1}{2\pi i}\int_{\gamma}R_{t}(\zeta)-R_{t_{0}}(\zeta)d\zeta\underset{t\to t_{0}}{=}\mathcal{O}(t-t_{0}),$
hence
$(I-D(t)^{2})^{-1/2}\underset{t\to t_{0}}{=}I+\mathcal{O}(|t-t_{0}|^{2}).$
Thus
$U(t)$ $=(I-D(t)^{2})^{-1/2}\left((I-P(t))(I-P(t_{0}))+P(t)P(t_{0})\right),$
$\hskip-5.69054pt\underset{t\to t_{0}}{=}\left((I-P(t))(I-P(t_{0}))+P(t)P(t_{0})\right)+o(t-t_{0}),$ (3.48)
$U(t)^{-1}$ $=\left((I-P(t_{0}))(I-P(t))+P(t_{0})P(t)\right)(I-D(t)^{2})^{-1/2},$
$\hskip-5.69054pt\underset{t\to t_{0}}{=}\left((I-P(t_{0}))(I-P(t))+P(t_{0})P(t)\right)+o(t-t_{0}).$ (3.49)
Using (3.48), (3.49) and (3.41) we find that
$P(t_{0})U(t)P(t_{0})\hskip-1.42262pt\underset{t\to t_{0}}{=}P(t_{0})P(t)P(t_{0})+o(t-t_{0})\hskip-1.42262pt\underset{t\to t_{0}}{=}P(t_{0})+o(t-t_{0}),$
$P(t_{0})U(t)^{-1}P(t_{0})\hskip-1.42262pt\underset{t\to t_{0}}{=}P(t_{0})P(t)P(t_{0})+o(t-t_{0})\hskip-1.42262pt\underset{t\to t_{0}}{=}P(t_{0})+o(t-t_{0}),$
as required. On the other hand, from (3.48) and (3.49) we have
$\left(I-P(t_{0})\right)U(t)P(t_{0})\hskip-1.42262pt\underset{t\to t_{0}}{=}\left(I-P(t_{0})\right)P(t)P(t_{0})+o(t-t_{0}),$
$P(t_{0})U(t)^{-1}\left(I-P(t_{0})\right)\hskip-1.42262pt\underset{t\to t_{0}}{=}P(t_{0})P(t)\left(I-P(t_{0})\right)+o(t-t_{0}).$
Thus, multiplying (3.42) on the left by $(I-P(t_{0}))$ yields (3.46), while multiplying (3.43) on the right by $(I-P(t_{0}))$ yields (3.47). ∎
Step 3. We have
$\begin{split}P(t_{0})&U^{-1}(t)R_{t}(\zeta)U(t)P(t_{0})\underset{t\to t_{0}}{=}(\lambda-\zeta)^{-1}P(t_{0})\\ &+(\lambda-\zeta)^{-2}\Big(-P(t_{0})\dot{V}_{t_{0}}P(t_{0})+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\left(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\right)\operatorname{T}_{t_{0}}P(t_{0})\\ &\qquad\qquad\qquad\qquad+\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}J\dot{\operatorname{T}}_{t_{0}}P(t_{0})\Big)(t-t_{0})+o(t-t_{0}).\end{split}$ (3.50)
Proof.
Sandwiching the factor $R_{t}(\zeta)$ in the left hand side by $P(t_{0})+(I-P(t_{0}))$, we have
$\begin{split}P(t_{0})U^{-1}(t)R_{t}(\zeta)U(t)P(t_{0})&=P(t_{0})U^{-1}(t)\Big(P(t_{0})+(I-P(t_{0}))\Big)R_{t}(\zeta)\\ &\qquad\qquad\qquad\qquad\times\Big(P(t_{0})+(I-P(t_{0}))\Big)U(t)P(t_{0}),\end{split}$
$\eqqcolon I+II+III+IV.$
We compute each term individually. We have
$I\coloneqq P(t_{0})U^{-1}(t)(I-P(t_{0}))\times(I-P(t_{0}))R_{t}(\zeta)P(t_{0})\times P(t_{0})U(t)P(t_{0})\underset{t\to t_{0}}{=}o(t-t_{0}),$
which follows from (3.39), (3.44) and (3.47), since the terms $(I-P(t_{0}))R_{t}(\zeta)P(t_{0})$ and $P(t_{0})U^{-1}(t)(I-P(t_{0}))$ are both $\mathcal{O}(t-t_{0})$. We similarly infer that
$II\coloneqq P(t_{0})U^{-1}(t)P(t_{0})\times P(t_{0})R_{t}(\zeta)(I-P(t_{0}))\times(I-P(t_{0}))U(t)P(t_{0})\underset{t\to t_{0}}{=}o(t-t_{0}),$
using (3.40), (3.45) and (3.46), and
$III\coloneqq P(t_{0})U^{-1}(t)(I-P(t_{0}))\times R_{t}(\zeta)\times(I-P(t_{0}))U(t)P(t_{0})\underset{t\to t_{0}}{=}o(t-t_{0})$
by (3.46) and (3.47). Finally, the term
$IV\coloneqq P(t_{0})U^{-1}(t)P(t_{0})\times P(t_{0})R_{t}(\zeta)P(t_{0})\times P(t_{0})U(t)P(t_{0})$ (3.51)
admits the sought after expansion (3.50), upon using (3.39), (3.38), (3.44) and (3.45). ∎
Step 4. Recalling the identities
$(\mathcal{N}_{t}+V_{t})P(t)\coloneqq\frac{-1}{2\pi i}\int_{\gamma}\zeta R_{t}(\zeta)d\zeta,\qquad\frac{1}{2\pi i}\int_{\gamma}\zeta(\lambda-\zeta)^{-1}d\zeta=1,$ (3.52)
multiplying (3.50) by $\zeta/2\pi i$ and integrating over $\gamma$, we arrive at (LABEL:operator_expansion). ∎
Theorem 3.10.
Assume Hypothesis 3.6 and that the mappings $t\mapsto\operatorname{T}_{t},t\mapsto V_{t},t\mapsto P_{t},t\mapsto Q_{t},$ are differentiable at $t_{0}$. We define
$T^{(1)}\coloneqq P(t_{0})\dot{V}_{t_{0}}P(t_{0})-\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\left(\dot{P}_{t_{0}}J\oplus-\dot{Q}_{t_{0}}J\right)\operatorname{T}_{t_{0}}P(t_{0})-\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0}),$
and denote the eigenvalue and eigenvector of this one-dimensional operator by $\lambda_{t_{0}}^{(1)}$ and $\{\mathbf{u}_{t_{0}}\}\subset\operatorname{ran}P(t_{0})=\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda)$, respectively. Suppose further that $\lambda\in\mathbb{R}$ then for $t$ near $t_{0}$, the eigenvalue $\lambda(t)$ of $\mathcal{N}_{t}+V_{t}$ satisfies the asymptotic formula
$\lambda(t)=\lambda+\lambda_{t_{0}}^{(1)}(t-t_{0})+o(t-t_{0}).$ (3.53)
Moreover, one has
$\lambda^{\prime}(t_{0})=\frac{\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right)}{\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}},$ (3.54)
where $\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}\neq 0$.
Proof.
Recalling that $U(t)$ is an isomorphism between $\operatorname{ran}P(t_{0})$ and $\operatorname{ran}P(t)$, see [Kato, Section I.4.6], [F04, Proposition 2.18], we note that $(\mathcal{N}_{t}+V_{t})|_{\operatorname{ran}P(t)}$ is similar to
$P(t_{0})U(t)^{-1}(\mathcal{N}_{t}+V_{t})P(t)U(t)P(t_{0})|_{\operatorname{ran}P(t_{0})}$
for $t$ near $t_{0}$. Thus the eigenvalues of these operators coincide for all $t$ near $t_{0}$, see [Kato, Section I.5.7]. Asymptotically expanding the eigenvalues of the latter using the finite dimensional first order perturbation theory, specifically [Kato, Theorem II.5.11], we deduce (3.53).
Next, applying $\langle\cdot,\tau\mathbf{u}_{t_{0}}\rangle$ to the eigenvalue equation $T^{(1)}\mathbf{u}_{t_{0}}=\lambda_{t_{0}}^{(1)}\mathbf{u}_{t_{0}}$, we find
$\lambda_{t_{0}}^{(1)}\langle\mathbf{u}_{t_{0}},\tau\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}=\left\langle P(t_{0})\dot{V}_{t_{0}}P(t_{0})\mathbf{u}_{t_{0}},\tau\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}$
$\quad-\left\langle\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\left(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}}\right)\mathcal{J}\operatorname{T}_{t_{0}}P(t_{0})\mathbf{u}_{t_{0}},\tau\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}-\left\langle\tau\left(\operatorname{T}_{t_{0}}P(t_{0})\right)^{*}\mathcal{J}\dot{\operatorname{T}}_{t_{0}}P(t_{0})\mathbf{u}_{t_{0}},\tau\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}},$
$\,\,=\left\langle\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},P(t_{0})^{*}\tau\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}-\left\langle\left(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}}\right)\mathcal{J}\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right\rangle_{\mathfrak{H}^{4}}-\left\langle\mathcal{J}\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right\rangle_{\mathfrak{H}^{4}}$
$\,\,=\left\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}+\left\langle\mathcal{J}\left(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}}\right)\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right\rangle_{\mathfrak{H}^{4}}-\Omega\left(\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right),$
$\,\,=\left\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}+\Omega\left(\left(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}}\right)\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right),$
where in the second last line, we used that $P(t_{0})^{*}\tau=\tau P(t_{0})$ , and that $\operatorname{ran}(P_{t}\oplus Q_{t})$ is Lagrangian, hence $\mathcal{J}(P_{t}\oplus Q_{t})+(P_{t}\oplus Q_{t})\mathcal{J}=\mathcal{J}$ and $\mathcal{J}(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})=-(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\mathcal{J}$. In the last line, we used that $\Omega\left(\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)\in\mathbb{R}$, which follows from differentiating $t\mapsto\Omega\left(\operatorname{T}_{t}\mathbf{u}_{t_{0}},\operatorname{T}_{t}\mathbf{u}_{t_{0}}\right)$ at $t_{0}$.
Since $\lambda^{\prime}(t_{0})=\lambda_{t_{0}}^{(1)}$, (3.54) follows. That $\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}\neq 0$ follows from the assumption that $\lambda$ is simple. Indeed, in this case, there are no generalised eigenvectors, i.e. no solution $\hat{\mathbf{u}}\in\operatorname{dom}(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda)$ to the inhomogeneous equation
$(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda)\hat{\mathbf{u}}=\mathbf{u}_{t_{0}}.$
Noting that $\ker(\mathcal{N}_{t_{0}}^{*}+V_{t_{0}}^{*}-\lambda)=\{\tau\mathbf{u}_{t_{0}}\}$ (since $\lambda\in\mathbb{R}$), the conclusion now follows from the Fredholm Alternative (which is applicable due to $\lambda$ being a point in the discrete spectrum of $\mathcal{N}_{t_{0}}+V_{t_{0}}$). ∎
4. Hadamard formulas via Lyaponuv-Schmidt
In this section, under certain assumptions we prove the existence of eigenvalue curves $\lambda(t)\in\operatorname{Spec}\left(\mathcal{N}_{t}+V_{t}\right)$ or $\lambda\in\operatorname{Spec}\left(\mathcal{N}_{t(\lambda)}+V_{t(\lambda)}\right)$, and compute local formulas for their derivatives. The analysis is similar to that in [CCLM23], however the situation herein differs due to the $t$-dependence of the Lagrangian planes describing the boundary conditions.
Throughout this section we assume Hypothesis 3.1 and the following.
Hypothesis 4.1.
Given $\lambda_{0}\in\mathbb{R}$ and $t_{0}\in(0,1)$, we assume:
  1. (i)
    $\lambda_{0}\notin\operatorname{Spec_{\rm ess}}(\mathcal{N}_{t_{0}}+V_{t_{0}})$.
  2. (ii)
    $\dim\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})=1$, with eigenfunction $\mathbf{u}_{t_{0}}$.
In addition, there exists an interval $\mathcal{I}\subset[0,1]$ containing $t_{0}$ and an integer $n\geq 1$ such that:
  • (iii)
    the mappings $t\mapsto V_{t},t\mapsto P_{t},t\mapsto Q_{t},t\mapsto\operatorname{T}_{t}$ are $C^{2}$ on $\mathcal{I}$, and
  • (iv)
    $\ker\big((N^{*}+V_{t}^{*}-\lambda)\tau\big)\cap\operatorname{dom}(\tau N)=\{0\}$ for all $t$ in $\mathcal{I}$.
The main results of this section are the following Hadamard formulas for the first derivatives of geometrically simple eigenvalues; the Maslov crossing forms appearing therein are defined in the next section.
Theorem 4.2.
Assume Hypothesis 4.1 and let $\mathbf{q}=\operatorname{T}_{t_{0}}\mathbf{u}$.
  1. (1)
    If $\mathfrak{m}_{\lambda_{0}}\neq 0$, then there is a $C^{1}$ curve $\lambda=\lambda(t)$ near $t_{0}$ such that
    $\lambda(t)\in\operatorname{Spec}\left(\mathcal{N}_{t}+V_{t}\right),\qquad\lambda^{\prime}(t_{0})=-\frac{\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})}{\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})}.$ (4.1)
  2. (2)
    If $\mathfrak{m}_{t_{0}}\neq 0$, then there is a $C^{1}$ curve $t=t(\lambda)$ near $\lambda_{0}$ such that
    $\lambda\in\operatorname{Spec}\left(\mathcal{N}_{t(\lambda)}+V_{t(\lambda)}\right),\qquad t^{\prime}(\lambda_{0})=-\frac{\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})}{\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})}.$ (4.2)
4.1. Maslov crossing forms
Towards the proof of Theorem 4.2, we begin by computing crossing forms. The following discussion mimics that in [LS20first, §4.5]. Recall from (3.3) that
$\operatorname{T}_{t}(\operatorname{dom}\mathcal{A}_{t}\times\operatorname{dom}\mathcal{B}_{t})=\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t},$ (4.3)
and also note that it follows from (2.10) that $\operatorname{dom}(N^{*}\tau)=\operatorname{dom}A^{*}\times\operatorname{dom}A^{*}$. For $\lambda\in\mathbb{R}$ and $t\in[0,1]$, we consider the following families of subspaces:
$\begin{split}\mathbb{K}_{\lambda,t}&\coloneqq\operatorname{T}_{t}\big(\ker\big((N^{*}+V_{t}^{*}-\lambda)\tau\big)\big)\subset\mathfrak{H}^{4},\\ \mathcal{F}_{t}&\coloneqq\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}\subset\mathfrak{H}^{4},\\ \Upsilon_{\lambda,t}&\coloneqq\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}\subset\mathfrak{H}^{4}\oplus\mathfrak{H}^{4},\\ \mathfrak{D}&\coloneqq\{\mathbf{q}=(q,q)^{\top}:q\in\mathfrak{H}^{4}\}\subset\mathfrak{H}^{4}\oplus\mathfrak{H}^{4}.\end{split}$ (4.4)
Let us note that
$\ker(\mathcal{N}_{t}+V_{t}-\lambda)\neq\{0\}\,\,\,\iff\,\,\,\mathbb{K}_{\lambda,t}\cap\mathcal{F}_{t}\neq\{0\}\,\,\,\iff\,\,\,\Upsilon_{\lambda,t}\cap\mathfrak{D}\neq\{0\}.$ (4.5)
In addition, as in the standard self-adjoint case the Cauchy data plane $\mathbb{K}_{\lambda,t}$ is Lagrangian for $\lambda\in\mathbb{R}$.
Lemma 4.3.
For all $\lambda\in\mathbb{R}$, the Cauchy data space $\mathbb{K}_{\lambda,t}$ is a Lagrangian subspace of $\mathfrak{H}^{4}$ with respect to the symplectic form $\Omega$ defined by $\mathcal{J}$ in (2.13).
Proof.
The result follows from the fact that the operator $(N^{*}+V_{t}^{*}-\lambda)\tau=(\tau N)^{*}+V_{t}^{*}\tau-\lambda\tau$ is a bounded perturbation of the adjoint $(\tau N)^{*}$ of a symmetric operator $\tau N$ satisfying the Green’s identity (2.21) with the trace operator $\operatorname{T}=\operatorname{T}_{t}$ and [BBvnkF98, Proposition 3.5]. ∎
To compute crossing forms, it will be convenient to view the Lagrangian plane $\Upsilon_{\lambda,t}$ as the graph of an infinitesimally symplectic operator, the regularity of which follows from previous results. Let $\Pi_{\lambda,t}$ be the orthogonal projection onto $\Upsilon_{\lambda,t}$ so that the mapping $t\mapsto\Pi_{\lambda,t}$ is continuously differentiable on $[0,1]$ for each $\lambda\in\mathbb{R}$, see e.g. [LS18, p. 480-481]. For $\lambda\in\mathbb{R}$ and $t_{0}\in[0,1]$ satisfying Hypothesis 4.1, there is an interval $\mathscr{I}\subseteq\mathcal{I}$ centred around $t_{0}$ and a $C^{1}$ family of operators $t\mapsto\mathscr{M}_{\lambda,t},t\in\mathscr{I}$ which is in $C^{1}(\mathscr{I},\mathcal{B}\left(\Upsilon_{\lambda,t_{0}},(\Upsilon_{\lambda,t_{0}})^{\top}\right))$ with $\mathscr{M}_{\lambda,t_{0}}=0$ such that
$\Upsilon_{\lambda,t}=\left\{\mathbf{q}+\mathscr{M}_{\lambda,t}\mathbf{q}:\mathbf{q}\in\Upsilon_{\lambda,t_{0}}\right\},\quad t\in\mathscr{I},$ (4.6)
see, e.g. [CJLS14, Lemma 3.8]. Then $(\lambda_{0},t_{0})$ is called a conjugate point if $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}\neq\{0\}$ (or either of the equivalent assertions in (4.5) hold). The $t$-crossing form for the Lagrangian plane $t\mapsto\Upsilon_{\lambda_{0},t}$ with respect to $\mathfrak{D}$ on the finite dimensional intersection $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}$ is defined to be
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{p})\coloneqq\frac{d}{dt}\big|_{t=t_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda_{0},t}\mathbf{p}),\qquad\mathbf{q},\mathbf{p}\in\Upsilon_{\lambda_{0},t_{0}},$ (4.7)
where $\tilde{\Omega}=\Omega\oplus(-\Omega)$ is a symplectic form on $\mathfrak{H}^{4}\oplus\mathfrak{H}^{4}$. Similarly, for the Lagrangian path $\lambda\mapsto\Upsilon_{\lambda,t_{0}}$, the $\lambda$-crossing form with respect to $\mathfrak{D}$ is defined as
$\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{p})\coloneqq\frac{d}{d\lambda}\big|_{\lambda=\lambda_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda,t_{0}}\mathbf{p}),\qquad\mathbf{q},\mathbf{p}\in\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}.$ (4.8)
Finally, following [PT09], when $\mathfrak{m}_{\lambda_{0}}=0$ we define the second order crossing form to be
$\mathfrak{m}^{(2)}_{\lambda_{0}}(\mathbf{q},\mathbf{p})=\frac{d^{2}}{d\lambda^{2}}\big|_{\lambda=\lambda_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda,t_{0}}\mathbf{p}).$ (4.9)
The following is an analogue of [LS20first, Lemma 4.19].
Lemma 4.4.
Suppose $(\lambda_{0},t_{0})$ is a conjugate point, i.e. $\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})\neq\{0\}$, or equivalently $\mathcal{K}_{\lambda_{0},t_{0}}\cap\mathcal{F}_{t_{0}}\neq\{0\}$. Let $\mathbf{u}_{t_{0}}\in\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})$. Then there exists an open interval $\mathscr{I}\subseteq\mathcal{I}$ around $t_{0}$, a family $t\mapsto\mathbf{w}_{t}$ in $C^{1}(\mathscr{I},\mathcal{H}_{+}^{2})$ and a family $t\mapsto g_{t}\in\mathcal{F}_{t}$ in $C^{1}(\mathscr{I},\mathfrak{H}^{4}\oplus\mathfrak{H}^{4})$ such that
$\mathbf{w}_{t_{0}}$ $=\mathbf{u}_{t_{0}},\quad g_{t_{0}}=\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},$ (4.10a)
$\mathbf{w}_{t}$ $\in\ker(N^{*}+V_{t}^{*}-\lambda_{0})\tau,$ (4.10b)
$(\operatorname{T}_{t}\mathbf{w}_{t},g_{t})^{\top}$ $=(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}})^{\top}+\mathscr{M}_{t,\lambda_{0}}(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}})^{\top}.$ (4.10c)
Similarly, under the same assumptions, there exists an open interval interval $\mathscr{I}\subseteq\mathcal{I}$ around $\lambda_{0}$ and a family $\lambda\mapsto\mathbf{w}_{\lambda}$ in $C^{\infty}(\mathscr{I},\mathcal{H}_{+}^{2})$ such that
$\mathbf{w}_{\lambda_{0}}$ $=\mathbf{u}_{t_{0}},$ (4.11a)
$\mathbf{w}_{\lambda}$ $\in\ker(N^{*}+V_{t_{0}}^{*}-\lambda)\tau,$ (4.11b)
$(\operatorname{T}_{t_{0}}\mathbf{w}_{\lambda},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}})^{\top}$ $=(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}})^{\top}+\mathscr{M}_{t_{0},\lambda}(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}})^{\top}$ (4.11c)
Proof.
We begin with the proof of (4.10) and the statements preceding those equations. The proof is almost identical to that of [LS20first, Lemma 4.19], but we give it here for completeness. Let us momentarily denote $K_{t}\coloneqq\ker\left((N^{*}+V_{t}-\lambda)\tau\right)$, so that $\mathbb{K}_{\lambda,t}=\operatorname{T}_{t}K_{t}$. We also denote $q\coloneqq\operatorname{T}_{t_{0}}\mathbf{u}$, $\mathbf{q}\coloneqq(q,q)$, and let $\mathbb{P}_{t}$ be the orthogonal projections onto $\mathbb{K}_{\lambda,t}$. Then $\mathbb{P}_{t}\in C^{1}(\mathcal{I},\mathcal{B}(\mathfrak{H}^{4}))$ for some open interval $\mathscr{I}\subseteq\mathcal{I}$ centred at $t_{0}$ (see, for example, [BBvnkF98, Theorem 3.9]).
We now consider the projections in $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ given by
$\widehat{\mathbb{P}}_{t}\coloneqq\begin{bmatrix}\mathbb{P}_{t}&0\\ 0&0\end{bmatrix},\qquad\widehat{\mathbb{Q}}_{t}\coloneqq\begin{bmatrix}0&0\\ 0&P_{t}\oplus Q_{t}\end{bmatrix}$, (4.12)
so that $\widehat{\mathbb{P}}_{t}+\widehat{\mathbb{Q}}_{t}=\Pi_{\lambda,t}$ where $\operatorname{ran}(\Pi_{\lambda,t})=\Upsilon_{\lambda,t}=\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}$. Using the definition of $\Upsilon_{\lambda,t}$ and $\mathscr{M}_{\lambda,t}$, see (4.4) and (4.6), we define
$h_{t}\in\operatorname{ran}\mathbb{P}_{t}\subset\mathfrak{H}^{4},\quad g_{t}\in\operatorname{ran}(P_{t}\oplus Q_{t})\subset\mathfrak{H}^{4}$ (4.13)
such that
$(h_{t},0)^{\top}=\widehat{\mathbb{P}}_{t}(\mathbf{q}+\mathscr{M}_{\lambda,t}\mathbf{q}),\qquad(0,g_{t})^{\top}=\widehat{\mathbb{Q}}(\mathbf{q}+\mathscr{M}_{\lambda,t}\mathbf{q}),$ (4.14)
and so $h_{t_{0}}=g_{t_{0}}=q$. Since $t\mapsto\mathscr{M}_{\lambda,t},t\mapsto\mathbb{P}_{t}$ and $t\mapsto P_{t}\oplus Q_{t}$ are $C^{1}$, it follows that $t\mapsto h_{t}$ and $t\mapsto g_{t}$ are $C^{1}$. Employing Hypothesis 4.1 and that $\ker(\operatorname{T}_{t})=\operatorname{dom}(N)$ (see (2.19)), it follows that the restriction
$\operatorname{T}_{t}|_{K_{t}}:K_{t}\to\operatorname{ran}\mathbb{P}_{t}\subset\mathfrak{H}^{4}$ (4.15)
is a bijection. Hence, there is a unique vector $\mathbf{w}_{t}\in K_{t}$ such that $\operatorname{T}_{t}\mathbf{w}_{t}=h_{t}$. Thus equations (4.10) hold for this choice of $\mathbf{w}_{t}$ and $g_{t}$.
It remains to show that the mapping $t\mapsto\mathbf{w}_{t}$ is in $C^{1}(\mathscr{I},\mathcal{H}_{+}^{2})$; we will exploit the bijectivity of the mapping (4.15), although some care is needed to handle the $t$-dependent domain of that mapping. Let $U_{t}$ denote the $C^{1}$ family of boundedly invertible transformation operators in $\mathcal{H}_{+}^{2}$ that split the projections $\mathcal{P}_{K_{t}}$ onto $K_{t}$ and $\mathcal{P}_{K_{t_{0}}}$ onto $K_{t_{0}}$ such that $U_{t}\mathcal{P}_{K_{t_{0}}}=\mathcal{P}_{K_{t}}U_{t}$ and $U_{t}:K_{t_{0}}\to K_{t}$ are bijections for $t$ near $t_{0}$, see [LSS18, Remark 2.4], [CJLS14, Remark 3.5], [Dalecki_Krein, Section IV.1], [F04, Remark 6.11]. Introducing $\mathbf{v}_{t}\in K_{t_{0}}$ by $\mathbf{v}_{t}=U_{t}^{-1}\mathbf{w}_{t}$ so that $\operatorname{T}_{t}\mathbf{w}_{t}=h_{t}$ yields $(\operatorname{T}_{t}\circ U_{t})\mathbf{v}_{t}=h_{t}$. The map $\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}:K_{t_{0}}\to\operatorname{ran}\mathbb{P}_{t}$ is a bijection and $t\mapsto\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}$ is in $C^{1}(\mathcal{I},\mathcal{B}(K_{t_{0}},\mathfrak{H}^{4}))$ by the assumptions in the lemma. Since $\mathbf{w}_{t}=U_{t}\circ(\operatorname{T}_{t}\circ U_{t})^{-1}h_{t}$, the function $t\mapsto\mathbf{w}_{t}$ is $C^{1}$ because each of the terms in the composition is $C^{1}$.
The proof above can now be applied to prove the existence and regularity of the family $\lambda\mapsto\mathbf{w}_{\lambda}$ satisfying (4.11). Indeed, setting $\tilde{V}^{*}_{\lambda}\coloneqq V^{*}_{t_{0}}-\lambda$, in this situation only the bounded perturbation $\tilde{V}^{*}_{\lambda}$ depends on $\lambda$, while the trace operator $\operatorname{T}_{t_{0}}$ and Lagrangian plane $\operatorname{ran}P_{t_{0}}\times\operatorname{ran}Q_{t_{0}}$ are $\lambda$-independent. The mapping $\lambda\mapsto\tilde{V}_{\lambda}^{*}$ is $C^{\infty}$, and the proof now follows from the arguments used to prove (4.10). ∎
We are ready to compute crossing forms.
Lemma 4.5.
Let $(\lambda_{0},t_{0})$ be a simple conjugate point, let $\mathbf{u}_{t_{0}}\in\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})$ and let $\mathbf{q}\coloneqq(\operatorname{T}\mathbf{u}_{t_{0}},\operatorname{T}\mathbf{u}_{t_{0}})\in\Upsilon_{\lambda,t}\cap\mathfrak{D}$. The $t$-crossing form introduced in (4.7) for the Lagrangian path $t\mapsto\Upsilon_{\lambda_{0},t}$ with respect to the reference plane $\mathfrak{D}$ (c.f. (4.4)) is given by
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})$ $=\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right)$ (4.16)
while the $\lambda$ crossing form introduced in (4.8) for the Lagrangian path $\lambda\mapsto\Upsilon_{\lambda,t_{0}}$ with respect to the reference plane $\mathfrak{D}$ is given by
$\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})$ $=-\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}.$ (4.17)
Proof.
We have
$(N^{*}+V_{t}^{*}-\lambda_{0})\tau\mathbf{w}_{t}=0,$ (4.18)
and, differentiating this equation with respect to $t$ at $t_{0}$ and applying $\langle\cdot,\mathbf{w}_{t_{0}}\rangle_{\mathcal{H}^{2}}$, we find that
$\langle(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau\dot{\mathbf{w}}_{t_{0}},\mathbf{w}_{t_{0}}\rangle_{\mathcal{H}^{2}}+\langle\dot{V}^{*}_{t_{0}}\tau\mathbf{w}_{t_{0}},\mathbf{w}_{t_{0}}\rangle_{\mathcal{H}^{2}}=0.$ (4.19)
Using Green’s identity (2.21) with $\mathbf{u}=\dot{\mathbf{w}}_{t_{0}}$ and $\mathbf{v}={\mathbf{w}}_{t_{0}}$, we find that
$\left\langle(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau\dot{\mathbf{w}}_{t_{0}},{\mathbf{w}}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}-\left\langle\dot{\mathbf{w}}_{t_{0}},(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau{\mathbf{w}}_{t_{0}}\right\rangle_{\mathcal{H}^{2}}=\Omega(\operatorname{T}_{t_{0}}\dot{\mathbf{w}}_{t_{0}},\operatorname{T}_{t_{0}}{\mathbf{w}}_{t_{0}}).$ (4.20)
Combining the previous two equations and noting that $\mathbf{w}_{t_{0}}\in\ker(N^{*}+V_{t_{0}}^{*}-\lambda_{0})\tau$, we obtain
$\langle\dot{V}^{*}_{t_{0}}\tau\mathbf{w}_{t_{0}},\mathbf{w}_{t_{0}}\rangle_{\mathcal{H}^{2}}-\Omega(\operatorname{T}_{t_{0}}{\mathbf{w}}_{t_{0}},\operatorname{T}_{t_{0}}\dot{\mathbf{w}}_{t_{0}})=0.$ (4.21)
Now by (4.7) and (4.10c) we have
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})$ $=\frac{d}{dt}\Big|_{t=t_{0}}\,\tilde{\Omega}((\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}},g_{t_{0}})^{\top},(\operatorname{T}_{t}\mathbf{w}_{t},g_{t})^{\top}),$
$=\Omega\left(\operatorname{T}\mathbf{w}_{t_{0}},\operatorname{T}\dot{\mathbf{w}}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{w}_{t_{0}}\right)-\Omega\left(g_{t_{0}},\dot{g}_{t_{0}}\right).$ (4.22)
Since $g_{t}\in\mathcal{F}_{t}$ and hence $g_{t}=(P_{t}\oplus Q_{t})g_{t}$, we have
$\dot{g}_{t_{0}}=(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})g_{t_{0}}+(P_{t_{0}}\oplus Q_{t_{0}})\dot{g}_{t_{0}}.$ (4.23)
Using this, along with the fact that $\mathcal{F}_{t}=\operatorname{ran}(P_{t})\times\operatorname{ran}(Q_{t})$ is a Lagrangian subspace of $\mathfrak{H}^{4}\oplus\mathfrak{H}^{4}$, and $\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}=\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}}\in\operatorname{ran}(P_{t_{0}})\times\operatorname{ran}(Q_{t_{0}})$, we find that
$\Omega\left(\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}},\dot{g}_{t_{0}}\right)$ $=\Omega\left(\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}},(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}}+(P_{t_{0}}\oplus Q_{t_{0}})\dot{g}_{t_{0}}\right),$
$=\Omega\left(\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}},(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{w}_{t_{0}}\right).$ (4.24)
Using (4.24) and (4.21) in (4.22), and the fact that $\dot{V}^{*}_{t_{0}}\tau=\tau\dot{V}_{t_{0}}$, yields (4.16).
For the $\lambda$-crossing form calculation, we now consider the family of vectors $\lambda\mapsto\mathbf{w}_{\lambda}$ given by (4.11). By (4.8) 333here we note that the second component of the mapping $\lambda\mapsto(\operatorname{T}_{t_{0}}\mathbf{w}_{\lambda},\operatorname{T}_{t_{0}}u_{t_{0}})^{\top}$ is $\lambda$-independent, hence, the computation reduces the first component $\Omega$ of the symplectic form $\tilde{\Omega}=\Omega\oplus(-\Omega)$ we have
$\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})$ $=\frac{d}{d\lambda}\Big|_{\lambda=\lambda_{0}}\,\Omega(\operatorname{T}_{t_{0}}\mathbf{w}_{\lambda_{0}},\operatorname{T}_{t_{0}}\mathbf{w}_{\lambda})=\Omega\left(\operatorname{T}_{t_{0}}\mathbf{w}_{\lambda_{0}},\operatorname{T}_{t_{0}}\dot{\mathbf{w}}_{\lambda_{0}}\right).$ (4.25)
As prior, for the equation
$(N^{*}+V^{*}_{t_{0}}-\lambda)\tau\mathbf{w}_{\lambda}=0,$ (4.26)
differentiating at $\lambda_{0}$ and applying $\langle\cdot,\mathbf{w}_{\lambda_{0}}\rangle_{\mathcal{H}^{2}}$ yields
$\langle(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau\dot{\mathbf{w}}_{\lambda_{0}},\mathbf{w}_{\lambda_{0}}\rangle_{\mathcal{H}^{2}}-\langle\tau\mathbf{w}_{\lambda_{0}},\mathbf{w}_{\lambda_{0}}\rangle_{\mathcal{H}^{2}}=0.$ (4.27)
Using Green’s identity (2.21) with $\mathbf{u}=\dot{\mathbf{w}}_{\lambda_{0}}$ and $\mathbf{v}={\mathbf{w}}_{\lambda_{0}}$, we find that
$\left\langle(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau\dot{\mathbf{w}}_{\lambda_{0}},{\mathbf{w}}_{\lambda_{0}}\right\rangle_{\mathcal{H}^{2}}-\left\langle\dot{\mathbf{w}}_{\lambda_{0}},(N^{*}+V^{*}_{t_{0}}-\lambda_{0})\tau{\mathbf{w}}_{\lambda_{0}}\right\rangle_{\mathcal{H}^{2}}=\Omega(\operatorname{T}_{t_{0}}\dot{\mathbf{w}}_{\lambda_{0}},\operatorname{T}_{t_{0}}{\mathbf{w}}_{\lambda_{0}}).$ (4.28)
Combining the previous two equations and noting that $\mathbf{w}_{\lambda_{0}}\in\ker(N^{*}+V_{t_{0}}^{*}-\lambda_{0})\tau$, we obtain
$\Omega(\operatorname{T}_{t_{0}}{\mathbf{w}}_{\lambda_{0}},\operatorname{T}_{t_{0}}\dot{\mathbf{w}}_{\lambda_{0}})=-\langle\tau\mathbf{w}_{t_{0}},\mathbf{w}_{t_{0}}\rangle_{\mathcal{H}^{2}}.$ (4.29)
Substituting the previous equation into the $\lambda$-crossing form, we obtain (4.17).
4.2. Lyaponuv-Schmidt reduction
In this subsection we prove the following proposition, which states that eigenvalues $\lambda\in\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$, for $(\lambda,t)$ near $(\lambda_{0},t_{0})$, are determined by the zero set of the determinant of a symmetric $n\times n$ matrix $M(\lambda,t)$.
Proposition 4.6.
Assume $\dim\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})=g$ with basis $\{\mathbf{u}_{t_{0}}^{(1)},\dots,\mathbf{u}_{t_{0}}^{(g)}\}$. There exists an $g\times g$ matrix $M(\lambda,t)$, defined near $(\lambda_{0},t_{0})$, such that $\lambda\in\operatorname{Spec}(\mathcal{N}_{t}+V_{t})$ if and only if $\det M(\lambda,t)=0$. This matrix satisfies $M(\lambda_{0},t_{0})=0$,
$\frac{\partial M_{ji}}{\partial t}(\lambda_{0},t_{0})$ $=\left<\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}}^{(i)},\mathbf{u}_{t_{0}}^{(j)}\right>+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}^{(j)}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}^{(j)}_{t_{0}}\right)$, (4.30)
$\frac{\partial M_{ji}}{\partial\lambda}(\lambda_{0},t_{0})$ $=-\left<\tau\mathbf{u}_{t_{0}}^{(i)},\mathbf{u}_{t_{0}}^{(j)}\right$>. (4.31)
Our goal is to construct a matrix $M(\lambda,t)$, the zero set of the determinant of which locally coincides with the real spectrum of $\mathcal{N}_{t}+V_{t}$ (i.e. for $(\lambda,t)$ near $(\lambda_{0},t_{0})$). We proceed with Lyaponuv-Schmidt reduction.
Proof.
The first step is to split the eigenvalue equation $(\mathcal{N}_{t}+V_{t}-\lambda)\mathbf{u}=0$ into two parts, one of which can always be solved uniquely. Let $\Pi$ denote the $\mathcal{H}$-orthogonal projection onto $\ker(\mathcal{N}^{*}_{t_{0}}+V^{*}_{t_{0}}-\lambda_{0})$, so that $I-\Pi$ is the projection onto $\ker(\mathcal{N}^{*}_{t_{0}}+V^{*}_{t_{0}}-\lambda_{0})^{\perp}=\operatorname{ran}(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})$.
It follows that $\lambda$ is an eigenvalue of $\mathcal{N}_{t}$ if and only if there exists a nonzero $\mathbf{u}\in\operatorname{dom}\mathcal{N}_{t}$ such that both
$\Pi(\mathcal{N}_{t}+V_{t}-\lambda)\mathbf{u}=0$ (4.32)
and
$(I-\Pi)(\mathcal{N}_{t}+V_{t}-\lambda)\mathbf{u}=0$ (4.33)
hold.
We first consider (4.33). Defining $X_{t}=\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})^{\perp}\cap\operatorname{dom}\mathcal{N}_{t}$, we have that any $\mathbf{u}\in\operatorname{dom}\mathcal{N}_{t}$ can be written uniquely as
$\mathbf{u}=\sum_{i=1}^{n}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}+\tilde{\mathbf{u}}$, (4.34)
where $\alpha_{i}\in\mathbb{R}$ and $\tilde{\mathbf{u}}\in X_{t}$. This means (4.33) holds if and only if there exists a vector $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{g})\in\mathbb{R}^{g}$ and a function $\tilde{\mathbf{u}}\in X_{t}$ such that
$(I-\Pi)(\mathcal{N}_{t}+V_{t}-\lambda)\left(\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}+\tilde{\mathbf{u}}\right)=0.$ (4.35)
We claim that for each $(\boldsymbol{\alpha},\lambda,t)$ there exists a unique $\tilde{\mathbf{u}}=\tilde{\mathbf{u}}(\boldsymbol{\alpha},\lambda,t)\in X_{t}$ satisfying (4.35). Writing this equation out explicitly, it is
$(I-\Pi)(\mathcal{N}_{t}+V_{t}-\lambda)\tilde{\mathbf{u}}(\boldsymbol{\alpha},\lambda,t)=-(I-\Pi)(\mathcal{N}_{t}+V_{t}-\lambda)\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}.$
We define
$T(\lambda,t)\colon X_{t}\to\operatorname{ran}(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0}),\qquad T(\lambda,t)=(I-\Pi)\big(\mathcal{N}_{t}+V_{t}-\lambda\big)\Big|_{X_{t}},$ (4.36)
and observe that $T(\lambda_{0},t_{0})$ is invertible, hence, due to continuity of resolvents as mappings from $\mathcal{H}^{2}$ to $\mathcal{H}^{2}_{+}$, see Proposition 3.5 (2), $T(\lambda,t)$ is also invertible for nearby $(\lambda,t)$.
In a slight abuse of notation, we denote $X_{t}^{\perp}=\ker(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})\cap\operatorname{dom}\mathcal{N}_{t}$. Then, defining
$W(\lambda,t):X_{t}^{\perp}\to X_{t},\qquad W(\lambda,t)=-T^{-1}(\lambda,t)(I-\Pi)\big(\mathcal{N}_{t}+V_{t}-\lambda\big)\Big|_{X_{t}^{\perp}},$ (4.37)
the unique solution to (4.35) is thus
$\tilde{\mathbf{u}}(\boldsymbol{\alpha},\lambda,t)=W(\lambda,t)\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}.$ (4.38)
So far we have shown that the equation $(I-\Pi)(\mathcal{N}_{t}+V_{t}-\lambda)\mathbf{u}=0$ is satisfied if and only if $\mathbf{u}$ has the form
$\mathbf{u}=\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}+W(\lambda,t)\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}=\big(I+W(\lambda,t)\big)\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}$ (4.39)
for some $\boldsymbol{\alpha}\in\mathbb{R}^{g}$. We conclude that there exists $\mathbf{u}$ for which $(\mathcal{N}_{t}+V_{t}-\lambda)\mathbf{u}=0$ holds if and only if
$\Pi(\mathcal{N}_{t}+V_{t}-\lambda)\big(I+W(\lambda,t)\big)\left(\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}\right)=0$ (4.40)
for some $\boldsymbol{\alpha}\in\mathbb{R}^{g}$. Moreover, $\mathbf{u}$ is nonzero if and only if $\boldsymbol{\alpha}$ is nonzero. Finally, we observe that $\ker(\mathcal{N}^{*}_{t_{0}}+V^{*}_{t_{0}}-\lambda_{0})$ is spanned by $\{\tau\mathbf{u}_{t_{0}}^{(1)},\tau\mathbf{u}_{t_{0}}^{(2)},\dots,\tau\mathbf{u}_{t_{0}}^{(g)}\}$, and so (4.40) is equivalent to
$\left<(\mathcal{N}_{t}+V_{t}-\lambda)\big(I+W(\lambda,t)\big)\left(\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}\right),\tau\mathbf{u}_{t_{0}}^{(j)}\right>=0,\qquad j=1,\dots,g.$ (4.41)
Defining the $g\times g$ matrix $M(\lambda,t)$ by
$M_{ji}(\lambda,t)=\left<(\mathcal{N}_{t}+V_{t}-\lambda)\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\tau\mathbf{u}_{t_{0}}^{(j)}\right>,\quad i,j=1,\dots,g,$ (4.42)
the system of $g$ equations (4.41) may be written as $M(\lambda,t)\boldsymbol{\alpha}=0$, which is satisfied for a nonzero vector $\boldsymbol{\alpha}$ if and only if $\det M(\lambda,t)=0$. This completes the first part of the proof.
It follows that $M(\lambda_{0},t_{0})=0$, because $W(\lambda_{0},t_{0})\mathbf{u}^{(i)}_{t_{0}}=0$. For the $t$ derivative, we first use the modified Green’s identity (2.20) and the fact that $\tau(\mathcal{N}_{t}+V_{t}-\lambda)$ is a symmetric operator to write
$M_{ji}(\lambda,t)$ $=\left<\tau(\mathcal{N}_{t}+V_{t}-\lambda)\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right$>,
$=\left<(\tau(N+V_{t_{0}}-\lambda))^{*}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right>+\left<(\tau V_{t}-\tau V_{t_{0}})^{*}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right$>,
$=\left<\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},(\tau(N+V_{t_{0}}-\lambda))^{*}\mathbf{u}_{t_{0}}^{(j)}\right>+\Omega\left(\operatorname{T}_{t}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\operatorname{T}_{t}\mathbf{u}^{(j)}_{t_{0}}\right)$
$\qquad\qquad\qquad+\left<(\tau V_{t}-\tau V_{t_{0}})^{*}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right$>,
$=\left<(\tau V_{t}-\tau V_{t_{0}})^{*}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right>+\Omega\left(\operatorname{T}_{t}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}},\operatorname{T}_{t}\mathbf{u}^{(j)}_{t_{0}}\right)$, (4.43)
because $(\tau(N+V_{t_{0}}-\lambda))^{*}\mathbf{u}_{t_{0}}^{(j)}=(\tau(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda))^{*}\mathbf{u}_{t_{0}}^{(j)}=0$. Now defining
$g_{t}\coloneqq\operatorname{T}_{t}\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}}$,
since $\big(I+W(\lambda,t)\big)\mathbf{u}^{(i)}_{t_{0}}\in\operatorname{dom}\mathcal{N}_{t}$ for all $t$ near $t_{0}$, we have $g_{t}\in\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}$. Hence $g_{t}=(P_{t}\oplus Q_{t})g_{t}$, and
$\dot{g}_{t_{0}}=(\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})g_{t_{0}}+({P}_{t_{0}}\oplus{Q}_{t_{0}})\dot{g}_{t_{0}}.$
Moreover, since $W(\lambda_{0},t_{0})\mathbf{u}^{(i)}_{t_{0}}=0$, we have $g_{t_{0}}=\operatorname{T}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}}$. Now differentiating (4.43) with respect to $t$ at $t_{0}$, we find that (where dot denotes $d/dt$),
$\frac{\partial M_{ji}}{\partial t}(\lambda_{0},t_{0})$ $=\left<(\tau\dot{V}_{t})^{*}\big(I+W(\lambda_{0},t)\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right>+\left<(\tau V_{t}-\tau V_{t_{0}})^{*}\big(I+\partial_{t}W(\lambda_{0},t_{0})\big)\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right>\Big|_{t=t_{0}}$
$\qquad\qquad\qquad+\Omega\left(\dot{g}_{t_{0}},\operatorname{T}_{t_{0}}\mathbf{u}^{(j)}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}^{(j)}_{t_{0}}\right)$,
$=\left<\tau\dot{V}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}},\mathbf{u}_{t_{0}}^{(j)}\right>+\Omega\left((\dot{P}_{t_{0}}\oplus\dot{Q}_{t_{0}})\operatorname{T}\mathbf{u}^{(i)}_{t_{0}},\operatorname{T}\mathbf{u}^{(j)}_{t_{0}}\right)+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}^{(i)}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}^{(j)}_{t_{0}}\right)$.
For the $\lambda$ derivative, we first observe that since $W(\lambda,t_{0})\mathbf{u}^{(i)}_{t_{0}}\in X_{t_{0}}\subseteq\operatorname{dom}\mathcal{N}_{t_{0}}=\operatorname{dom}\mathcal{A}_{t_{0}}\times\operatorname{dom}\mathcal{B}_{t_{0}}$ for all $\lambda$ near $\lambda_{0}$, we have $\partial_{\lambda}W(\lambda,t_{0})\mathbf{u}^{(i)}_{t_{0}}\in\operatorname{dom}\mathcal{N}_{t_{0}}$ for all $\lambda$ near $\lambda_{0}$, and hence
$\operatorname{T}_{t_{0}}\partial_{\lambda}W(\lambda_{0},t_{0})\mathbf{u}^{(i)}_{t_{0}}\in\operatorname{ran}P_{t_{0}}\times\operatorname{ran}Q_{t_{0}}.$ (4.44)
Therefore, we may differentiate (4.42) directly because $\partial_{\lambda}W(\lambda,t_{0})\mathbf{u}^{(i)}_{t_{0}}\in\operatorname{dom}\mathcal{N}_{t_{0}}$, arriving at
$\frac{\partial M_{ji}}{\partial\lambda}(\lambda_{0},t_{0})$ $=\left<-\big(I+W(\lambda_{0},t_{0})\big)\mathbf{u}^{(i)}_{t_{0}}+(\mathcal{N}_{t_{0}}+V_{t_{0}}-\lambda_{0})\partial_{\lambda}W(\lambda_{0},t_{0})\mathbf{u}^{(i)}_{t_{0}},\tau\mathbf{u}^{(j)}_{t_{0}}\right$>,
$=-\left<\mathbf{u}^{(i)}_{t_{0}},\tau\mathbf{u}^{(j)}_{t_{0}}\right$>,
where we used that $\tau\mathbf{u}^{(j)}_{t_{0}}\in\ker(\mathcal{N}_{t_{0}}^{*}+V_{t_{0}}^{*}-\lambda_{0})$, as required.
Proof of Theorem 4.2.
Suppose that $(\lambda_{0},t_{0})$ is a simple conjugate point. In this case $M(\lambda,t)$ defined in (4.42) is a scalar, and comparing the expressions in Lemma 4.5 and Proposition 4.6, we have
$\frac{\partial M}{\partial t}(\lambda_{0},t_{0})=\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q}),\qquad\frac{\partial M}{\partial\lambda}(\lambda_{0},t_{0})=\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q}),$ (4.45)
where $\mathbf{q}=\operatorname{T}\mathbf{u}_{t_{0}}$. Theorem 4.2 now follows from Proposition 4.6 and the implicit function theorem. Namely, in the case that $\partial_{\lambda}M(\lambda_{0},t_{0})=\mathfrak{m}_{\lambda_{0}}(\mathbf{q},\mathbf{q})\neq 0$, applying the implicit function theorem to $M(\lambda,t)=0$ implies that the existence of a $C^{1}$ curve $\lambda(t)$, defined for $t$ near $t_{0}$, whose first derivative $\lambda^{\prime}(t_{0})$ is given by (4.1). Formula (4.2) in the case when $\partial_{t}M(\lambda_{0},t_{0})=\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})\neq 0$ follows similarly. ∎
5. Application: Linearised NLS for standing waves on compact star graphs
In this section we apply our abstract theory to study the example detailed in the introduction, that is, linearisation about a standing wave solution (1.3) to the nonlinear Schrödinger equation (1.2) on a compact star graph $\mathcal{G}$, which satisfies the standing wave equation (1.4) and vertex conditions (1.5). After restricting the eigenvalue problem (1.7a)–(1.7b) to the sub-graph $\mathcal{G}_{t}$, $t\in(0,1]$, and rescaling back to $\mathcal{G}$, we obtain the $t$-dependent eigenvalue problem (1.10)–(1.11).
Our primary goal will be to prove Theorem 1.2. This follows from a homotopy argument and explicit expressions for the crossing forms, which are used to compute local contributions to the Maslov index, see (5.11). Our first task will therefore be to compute the crossing forms (4.7) and (4.8) in $t$ and $\lambda$ respectively. As an aside, we also compute Hadamard formulas for the first derivatives of the eigenvalue curves. We emphasise that the Hadamard-type formula for arbitrary $(\lambda,t)=(\lambda_{0},t_{0})$, given in (5.10), is inconsequential to the proof of Theorem 1.2; however, for the purposes of this paper, we write down this expression to highlight examples of the abstract Hadamard formulas given in Theorems 4.2 and 3.10.
Remark 5.1.
We will not discuss here the issue of the existence of a solution to (1.4) – (1.5). Since the focus will be on the spectrum of the associated linearised operator, instead we assume a solution exists, and use the crossing forms and Hadamard formulas derived in Section 4 to study the existence of positive real eigenvalues.
The $t$-dependent eigenvalue problem (1.10) is described in the notation of Section 2 as follows. The function spaces are
$\mathcal{H}=L^{2}(\mathcal{G}),\quad\mathcal{H}_{+}=\widehat{H}^{2}(\mathcal{G}),\quad\mathfrak{H}=L^{2}(\partial\mathcal{G})\cong\mathbb{R}^{2m}.$ (5.1)
The Sobolev space of functions vanishing on the boundary $\partial\mathcal{G}$ together with their derivatives is denoted by
$\widehat{H}^{2}_{0}(\mathcal{G})\coloneqq\left\{f\in\widehat{H}^{2}(\mathcal{G}):\operatorname{T}f=0\right\}.$
The minimal symmetric operator with finite and equal deficiency indices acting in $\mathcal{H}^{2}=\left(L^{2}(\mathcal{G})\right)^{2}$ is then $A=-\partial_{xx},\,\operatorname{dom}(A)=\widehat{H}_{0}^{2}(\mathcal{G})$, with maximal adjoint operator $A^{*}=-\partial_{xx},\,\operatorname{dom}(A^{*})=\mathcal{H}_{+}=\widehat{H}^{2}(\mathcal{G})$. The trace operators
$\operatorname{tr}_{t}=(\Gamma_{0,t},\Gamma_{1,t})^{\top}:\mathcal{H}_{+}\to\mathfrak{H},\qquad\operatorname{T}_{t}=\operatorname{tr}_{t}\oplus\operatorname{tr}_{t}=[\Gamma_{0,t},\Gamma_{1,t}]^{\top}\oplus[\Gamma_{0,t},\Gamma_{1,t}]^{\top}:\mathcal{H}_{+}^{2}\to\mathfrak{H}^{4},$
from Hypothesis 3.1 are given by (1.12), (1.9). Recalling that $\mathcal{L}\subset\mathfrak{H}^{4}$ is the Lagrangian plane describing the vertex conditions (1.5), we denote by $P_{\mathcal{L}}$ the ($t$-independent) orthogonal projection in $\mathbb{R}^{4m}$ onto $\mathcal{L}$. In the notation of Hypothesis 3.1, we therefore have $P_{t}=Q_{t}=P_{\mathcal{L}}$. Thus, the vertex conditions (1.5) state that $\operatorname{T}_{t}\mathbf{u}\in\mathcal{L}\oplus\mathcal{L}$ for $\mathbf{u}=(u,v)^{\top}\in\mathcal{H}_{+}^{2}$ if
$\operatorname{T}_{t}\mathbf{u}\in\mathcal{L}\oplus\mathcal{L}\implies\begin{cases}u_{1}(0)=u_{2}(0)=\dots=u_{m}(0),\\ u_{1}(\ell_{1})=u_{2}(\ell_{2})=\dots=u_{m}(\ell_{m})=0,\\ \sum_{i=1}^{m}u_{i}^{\prime}(0)=\alpha u_{1}(0),\quad\alpha\in\mathbb{R},\end{cases}$ (5.2)
with similar conditions holding for $v$.
Having established the relevant notation, we proceed with the computation of crossing forms. In what follows, we denote the spectral parameter in the eigenvalue problem (1.10) by $\mu\coloneqq t^{2}\lambda$. Let $(\lambda_{0},t_{0})$ be a conjugate point with eigenfunction $\mathbf{u}_{t_{0}}$. For the Lagrangian path $t\mapsto\Upsilon_{\mu_{0},t}$, the crossing form with respect to the diagonal subspace $\mathfrak{D}$ introduced in (4.7) is given by
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})$ $=\big\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\big\rangle_{\mathcal{H}^{2}}+\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right).$ (5.3)
For the first term in (5.3), using the expression for $V_{t}$ given above, and writing $\mathbf{u}_{t_{0}}=(u_{t_{0}},v_{t_{0}})\in\mathcal{H}^{2}$, we have
$\big\langle\tau\dot{V}_{t_{0}}$ $\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\big\rangle_{\mathcal{H}^{2}}=-\big\langle(2t_{0}g(t_{0}x)+t_{0}^{2}g^{\prime}(t_{0}x)x)u_{t_{0}},u_{t_{0}}\big\rangle_{\mathcal{H}}+\big\langle(2t_{0}h(t_{0}x)+t_{0}^{2}h^{\prime}(t_{0}x)x)v_{t_{0}},v_{t_{0}}\big\rangle_{\mathcal{H}},$
$=-\sum_{i=1}^{m}\int_{0}^{\ell_{i}}(2t_{0}g(t_{0}x)+t_{0}^{2}g^{\prime}(t_{0}x)x)u_{t_{0},i}^{2}(x)dx+\sum_{i=1}^{m}\int_{0}^{\ell_{i}}(2t_{0}h(t_{0}x)+t_{0}^{2}h^{\prime}(t_{0}x)x)v^{2}_{t_{0}}(x)dx.$
A direct calculation using the equation
$v_{t_{0}}^{\prime\prime}(x)+t_{0}^{2}h(t_{0}x)v_{t_{0}}=0$
shows that, for each $i=1,\dots,m$, we have
$\frac{d}{dx}\left[\frac{1}{t_{0}^{2}}x(v^{\prime}_{t_{0},i}(x))^{2}+x(v_{t_{0},i}(x))^{2}h(t_{0}x)-\frac{1}{t_{0}^{2}}v_{t_{0},i}(x)v^{\prime}_{t_{0},i}(x)\right]$ $=\left[2h(t_{0}x)+t_{0}xh^{\prime}(t_{0}x)\right]v_{t_{0},i}^{2}(x).$
Hence
$\int_{0}^{\ell_{i}}(2t_{0}h(t_{0}x)+t_{0}^{2}h^{\prime}(t_{0}x)x)v_{t_{0},i}^{2}(x)dx$ $=\frac{\ell_{i}}{t_{0}^{2}}(v^{\prime}_{t_{0},i}(\ell_{i}))^{2}+\frac{1}{t_{0}^{2}}v_{t_{0},i}(0)v_{t_{0},i}^{\prime}(0).$
It can be similarly deduced using $u_{t_{0}}^{\prime\prime}(x)+t_{0}^{2}g(t_{0}x)u_{t_{0}}=0$ that
$\int_{0}^{\ell_{i}}(2t_{0}g(t_{0}x)+t_{0}^{2}g^{\prime}(t_{0}x)x)u_{t_{0},i}^{2}(x)dx=\frac{\ell_{i}}{t_{0}^{2}}(u^{\prime}_{t_{0},i}(\ell_{i}))^{2}+\frac{1}{t_{0}^{2}}u_{t_{0},i}(0)u_{t_{0},i}^{\prime}(0),$
and therefore
$\big\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\big\rangle_{\mathcal{H}^{2}}$ $=-\sum_{i=1}^{m}\left(\frac{\ell_{i}}{t_{0}^{2}}(u^{\prime}_{t_{0},i}(\ell_{i}))^{2}+\frac{1}{t_{0}^{2}}u_{t_{0},i}(\ell_{i})u_{t_{0},i}^{\prime}(0),\right)+\sum_{i=1}^{m}\left(\frac{\ell_{i}}{t_{0}^{2}}(v^{\prime}_{t_{0},i}(\ell_{i}))^{2}+\frac{1}{t_{0}^{2}}v_{t_{0},i}(0)v_{t_{0},i}^{\prime}(0)\right).$
Using the conditions at the central vertex (the third line in (5.2) and the equivalent statement for $v_{t_{0}}$), we conclude that
$\big\langle\tau\dot{V}_{t_{0}}\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\big\rangle_{\mathcal{H}^{2}}$ $=\frac{1}{t_{0}^{2}}\sum_{i=1}^{m}\left\{\ell_{i}v^{\prime}_{t_{0},i}(\ell_{i})^{2}-\ell_{i}u^{\prime}_{t_{0},i}(\ell_{i})^{2}-u_{t_{0},i}(0)u_{t_{0},i}^{\prime}(0)+v_{t_{0},i}(0)v_{t_{0},i}^{\prime}(0)\right\},$
$=\frac{1}{t_{0}^{2}}\sum_{i=1}^{m}\left\{\ell_{i}v^{\prime}_{t_{0},i}(\ell_{i})^{2}-\ell_{i}u^{\prime}_{t_{0},i}(\ell_{i})^{2}\right\}-\frac{\alpha}{t_{0}^{2}}\left(u_{t_{0},1}^{2}(0)-v_{t_{0},1}^{2}(0)\right).$ (5.4)
For the second term in (5.3), from the definition of $\operatorname{T}_{t}$, we have
$\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}=(0,\dot{\Gamma}_{1,t_{0}}u_{t_{0}},0,\dot{\Gamma}_{1,t_{0}}v_{t_{0}})^{\top},$
where, for example,
$\dot{\Gamma}_{1,t_{0}}u_{t_{0}}=-\frac{1}{t_{0}^{2}}\left(u_{t_{0},1}^{\prime}(0),\dots,u_{t_{0},m}^{\prime}(0),-u_{t_{0},1}^{\prime}(\ell_{1}),\dots,-u_{t_{0},m}^{\prime}(\ell_{m})\right).$
Hence, from the definition of $\Omega$ in (2.13), and with $\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}}=(\Gamma_{0,t_{0}}u_{t_{0}},\Gamma_{1,t_{0}}u_{t_{0}},\Gamma_{0,t_{0}}v_{t_{0}},\Gamma_{1,t_{0}}v_{t_{0}})^{\top}$, we have
$\Omega\left(\operatorname{T}_{t_{0}}\mathbf{u}_{t_{0}},\dot{\operatorname{T}}_{t_{0}}\mathbf{u}_{t_{0}}\right)$ $=-\langle\Gamma_{0,t}u_{t_{0}},\dot{\Gamma}_{1,t}u_{t_{0}}\rangle_{\mathbb{R}^{2m}}+\langle\Gamma_{0,t}v_{t_{0}},\dot{\Gamma}_{1,t}v_{t_{0}}\rangle_{\mathbb{R}^{2m}}.$ (5.5)
Again using the vertex condition at the central vertex in (5.2), we find
$\langle\Gamma_{0,t}u_{t_{0}},\dot{\Gamma}_{1,t}u_{t_{0}}\rangle_{\mathbb{R}^{2m}}$ $=-\frac{1}{t_{0}^{2}}\sum_{i=1}^{m}u_{t_{0},i}(0)u^{\prime}_{t_{0},i}(0)=-\frac{u_{t_{0},1}(0)}{t_{0}^{2}}\sum_{i=1}^{m}u^{\prime}_{t_{0},i}(0)=-\frac{\alpha}{t_{0}^{2}}\left(u_{t_{0},1}(0)\right)^{2}.$ (5.6)
Similarly,
$\langle\Gamma_{0t}v_{t_{0}},\dot{\Gamma}_{1t}v_{t_{0}}\rangle_{\mathbb{R}^{2m}}=-\frac{\alpha}{t_{0}^{2}}\left(v_{t_{0},1}(0)\right)^{2}.$ (5.7)
Combining equations (5.4) and (5.5)–(5.7), (5.3) reduces to
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{q})$ $=\frac{1}{t_{0}^{2}}\sum_{i=1}^{m}\ell_{i}\left(v^{\prime}_{t_{0},i}(\ell_{i})^{2}-u^{\prime}_{t_{0},i}(\ell_{i})^{2}\right).$ (5.8)
On the other hand, the crossing form for the path $\mu\mapsto\Upsilon_{\mu,t_{0}}$ with respect to $\mathfrak{D}$ is given by
$\mathfrak{m}_{\mu_{0}}(\mathbf{q},\mathbf{q})$ $=-\langle\tau\mathbf{u}_{t_{0}},\mathbf{u}_{t_{0}}\rangle_{\mathcal{H}^{2}}=-2\langle u_{t_{0}},v_{t_{0}}\rangle_{\mathcal{H}}.$ (5.9)
Next, we use Theorem 4.2 to write down Hadamard-type variational formulas for the eigenvalue curves via crossing forms. Under the assumption that $\mu\in\operatorname{Spec}(\mathcal{N}_{t_{0}}+V_{t_{0}})$ is simple, we have $\mathfrak{m}_{\mu_{0}}\neq 0$. By Theorem 4.2 (or indeed Theorem 3.10) it follows that there exists a locally defined $C^{1}$ curve $\mu=\mu(t)$ satisfying $\mu(t_{0})=\mu_{0}=t_{0}^{2}\lambda_{0}$ and such that $\mu^{\prime}(t_{0})=-\mathfrak{m}_{t_{0}}/\mathfrak{m}_{\mu_{0}}$. Since $\mu=t^{2}\lambda$, the existence of a $C^{1}$ curve $\lambda(t)$ through $(t_{0}^{2}\lambda_{0},t_{0})$ immediately follows. By the chain rule we have $\mu^{\prime}(t)=\frac{d}{dt}(t^{2}\lambda(t))=2t\lambda(t)+t^{2}\lambda^{\prime}(t)$; rearranging yields the desired Hadamard formula for $\lambda^{\prime}(t_{0})$,
$\lambda^{\prime}(t_{0})=\frac{\sum_{i=1}^{m}\ell_{i}\left(v^{\prime}_{t_{0},i}(\ell_{i})^{2}-u^{\prime}_{t_{0},i}(\ell_{i})^{2}\right)-4\,t_{0}\lambda_{0}\,\langle u_{t_{0}},v_{t_{0}}\rangle_{\mathcal{H}}}{2\,t_{0}^{2}\,\langle u_{t_{0}},v_{t_{0}}\rangle_{\mathcal{H}}}.$ (5.10)
In the case when $\mathfrak{m}_{t_{0}}\neq 0$, by Theorem 4.2 there exists a locally defined $C^{1}$ curve $t(\lambda)$ through $(\lambda_{0},t_{0})$, whose derivative $t^{\prime}(\lambda_{0})$ is given by the reciprocal of the right hand side of (5.10).
We conclude with the proof of Theorem 1.2. We first recall a definition of the Maslov index via crossing forms relevant for our analysis, following [RS93]. Let $s$ denote a general parameter, representing either $t$ or $\lambda$, and denote by $s\mapsto\Upsilon_{s}$ the path obtained from the Lagrangian subspace $\Upsilon_{\lambda,t}$ by varying one of $t$ or $\lambda$, and holding the other fixed. For the Lagrangian path $s\mapsto\Upsilon_{s}$, let $s_{0}$ be a crossing, i.e. $\Upsilon_{s_{0}}\cap\mathfrak{D}\neq\{0\}$, and suppose $s_{0}\in[a,b]$ for some interval $[a,b]$ such that $s_{0}$ is the only crossing in $[a,b]$. Supposing that $s_{0}$ is regular, i.e. $\mathfrak{m}_{s_{0}}$ is nondegenerate, the Maslov index is defined locally as follows,
$\operatorname{Mas}(\Upsilon_{s},\mathfrak{D}:s\in[a,b])\coloneqq\begin{cases}-n_{-}(\mathfrak{m}_{s_{0}})&s_{0}=a,\\ n_{+}(\mathfrak{m}_{s_{0}})-n_{-}(\mathfrak{m}_{t_{0}})&a<s_{0}<b,\\ n_{+}(\mathfrak{m}_{s_{0}})&s_{0}=b,\end{cases}$ (5.11)
where $n_{+}(\mathfrak{m}_{s_{0}})$, resp. $n_{-}(\mathfrak{m}_{s_{0}})$, is the number of positive, resp. negative, squares of the quadratic form $\mathfrak{m}_{s_{0}}$. The Maslov index of the path $\mathscr{J}\ni s\mapsto\Upsilon_{s}$ is then obtained by summing the Maslov indices of each crossing $s_{0}\in\mathscr{J}$. Importantly, the Maslov index is invariant under fixed-endpoint homotopies of the path, and using this property one can extend the definition to all Lagrangian paths (i.e. those with non-regular crossings).
Next, we give an outline of the proof and record some preliminary results. Recall that $\mathbb{K}_{\lambda,t}$ is the Cauchy data plane, defined in (4.4), and $\mathcal{L}\subset\mathfrak{H}^{2}\times\mathfrak{H}^{2}$ is the Lagrangian plane describing the vertex conditions (5.2), i.e. $\delta$-type conditions at the central vertex and Dirichlet conditions at the free vertices, and $\mathfrak{D}\subset\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ is the diagonal plane in $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$. In what follows, we will exploit (4.5), i.e. that
$\ker(\mathcal{N}+V_{t}-t^{2}\lambda)\neq\{0\}\,\,\,\iff\,\,\,\Upsilon_{\lambda,t}\cap\mathfrak{D}\neq\{0\}$ (5.12)
(recalling that $t^{2}\lambda$ is the spectral parameter). For example, to show the triviality of intersections of Lagrangian planes, we will instead prove the triviality of the kernel of the associated differential operator.
Consider the Lagrangian path
$\Gamma\ni(\lambda,t)\mapsto\Upsilon_{\lambda,t}\coloneqq\mathbb{K}_{\lambda,t}\oplus(\mathcal{L}\oplus\mathcal{L})$ (5.13)
over the contour $\Gamma\coloneqq\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\cup\Gamma_{4}$ in the $\lambda t$-plane, oriented clockwise, where the segments $\Gamma_{i}$ are given by
$\begin{split}&\Gamma_{1}:\lambda\in[0,\lambda_{\infty}],\quad t=\varepsilon_{0};\qquad\Gamma_{2}:\lambda=0,\quad\,\,\,\,\,t\in[\varepsilon_{0},1];\\ &\Gamma_{3}:\lambda\in[0,\lambda_{\infty}],\quad t=1;\qquad\hskip 3.98337pt\Gamma_{4}:\lambda=\lambda_{\infty},\quad t\in[\varepsilon_{0},1].\end{split}$ (5.14)
Here $0<\varepsilon_{0}\ll 1$ and $\lambda_{\infty}$ is taken large enough so that
$\operatorname{Spec}(\mathcal{N}_{t}+V_{t})\cap\{z\in\mathbb{C}:{\rm{Re}}\,z\geq\lambda_{\infty}\}=\emptyset\quad\text{ for all}\quad t\in[\varepsilon_{0},1].$ (5.15)
Indeed such a choice is possible, since for each $t\in[\varepsilon_{0},1]$, $\mathcal{N}_{t}+V_{t}$ is a bounded perturbation of a skew self-adjoint operator, and therefore its spectrum is contained in the vertical strip $\left\{z\in\mathbb{C}:|{\rm{Re}}(z)|\leq\|V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}\right\}$ around the imaginary axis in the complex plane [Kato]. Taking the supremum of the spectral bounds over $t\in[\varepsilon_{0},1]$ yields the claim.
If $\varepsilon_{0}>0$ is small enough, we claim that $\operatorname{Spec}(\mathcal{N}+V_{\varepsilon_{0}})\cap\mathbb{R}=\emptyset$; the proof is similar to that of [CCLM23, Lemma 3.23]. First, we note that the operators $\mathcal{A}+F_{\varepsilon_{0}}$ and $\mathcal{A}+G_{\varepsilon_{0}}$ with domain $\operatorname{dom}(\mathcal{A})=\{u\in\mathcal{H}_{+}:\operatorname{tr}_{t}u\in\mathcal{L}\}$ are strictly positive. To see this, since $\alpha>0$ we note that $\operatorname{Spec}(\mathcal{A})\subset(0,\infty)$, hence, $\mathcal{A}\geq C$ for sufficiently small $C>0$. Using this inequality we arrive at
$\left\langle(\mathcal{A}+G_{\varepsilon_{0}})u,u\right\rangle_{L^{2}(\mathcal{G})}\geq C\|u\|^{2}_{L^{2}(\mathcal{G})}-\varepsilon_{0}^{2}\|g\|_{L^{\infty}(\mathcal{G})}\|u\|^{2}_{L^{2}(\mathcal{G})}>c\|u\|^{2}_{L^{2}(\mathcal{G})}$ (5.16)
for sufficiently small $c>0$, where $g(x)=(2p+1)\phi(x)+\beta$. A similar inequality holds for $\mathcal{A}+F_{\varepsilon_{0}}$. Next, suppose, by way of contradiction, that there exists $\varepsilon_{0}^{2}\lambda\in\operatorname{Spec}(\mathcal{N}_{\varepsilon_{0}})\cap\mathbb{R}$ with eigenfunction $\mathbf{u}_{t_{0}}=(u_{t_{0}},v_{t_{0}})\in\operatorname{dom}(\mathcal{N}_{\varepsilon_{0}})$. Using the fact that $\mathcal{A}+G_{\varepsilon_{0}}$ is strictly positive and arguing as in [CCLM23, Lemma 3.21] we note that the eigenvalue problem (1.10) for $t=\varepsilon_{0}$ small enough is equivalent to the self-adjoint eigenvalue problem
$(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}(\mathcal{A}+G_{\varepsilon_{0}})(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}w_{\varepsilon_{0}}=-\varepsilon_{0}^{4}\lambda^{2}w_{\varepsilon_{0}},$ (5.17)
where $w_{\varepsilon_{0}}\in\operatorname{dom}(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}$ is such that $(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}w_{\varepsilon_{0}}\in\operatorname{dom}(\mathcal{A}+G_{\varepsilon_{0}})$ and $(\mathcal{A}+G_{\varepsilon_{0}})(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}w_{\varepsilon_{0}}\in\operatorname{dom}(\mathcal{A}+F_{\varepsilon_{0}})$. Applying $\langle\cdot,w_{\varepsilon_{0}}\rangle$ to (5.17), using self-adjointness of $(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}$ and positivity of $\mathcal{A}+F_{\varepsilon_{0}},\mathcal{A}+G_{\varepsilon_{0}}$, for positive constants $C_{G},C_{F}$ we find that
$0>-\varepsilon_{0}^{4}\lambda^{2}\|w_{\varepsilon_{0}}\|^{2}=\langle(\mathcal{A}+G_{\varepsilon_{0}})(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}w_{\varepsilon_{0}},(\mathcal{A}+F_{\varepsilon_{0}})^{1/2}w_{\varepsilon_{0}}\rangle\geq C_{G}C_{F}\|w_{\varepsilon_{0}}\|^{2}>0.$ (5.18)
We conclude that
$\operatorname{Spec}(\mathcal{N}+V_{\varepsilon_{0}})\cap\mathbb{R}=\emptyset.$ (5.19)
Proof of Theorem 1.2.
By homotopy invariance and additivity under concatenation we have
$\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{1})+\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{2})+\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{3})+\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{4})=0.$ (5.20)
From the choice of $\lambda_{\infty}$ given by (5.15), it follows that $\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{4})=0$. In addition, for our choice of $\varepsilon_{0}>0$ small, by (5.19) we have $\operatorname{Mas}(\Upsilon_{\lambda,t},\mathfrak{D};\Gamma_{1})=0$. Again using the concatenation property, it follows that
$\operatorname{Mas}(\Upsilon_{0,t},\mathfrak{D};t\in[\varepsilon_{0},1-\varepsilon])+\mathfrak{c}+\operatorname{Mas}(\Upsilon_{\lambda,1},\mathfrak{D};\lambda\in[\varepsilon,\lambda_{\infty}])=0,$ (5.21)
for some $\varepsilon>0$, where
$\mathfrak{c}\coloneqq\operatorname{Mas}(\Upsilon_{0,t},\mathfrak{D};t\in[1-\varepsilon,1])+\operatorname{Mas}(\Upsilon_{\lambda,1},\mathcal{D};\lambda\in[0,\varepsilon]).$ (5.22)
By choosing $\varepsilon>0$ small enough, we can guarantee that $\mathfrak{c}$ represents the contribution to the Maslov index from the corner crossing $(\lambda,t)=(0,1)$ only. This follows from the fact that $(\lambda,t)=(0,1)$ is an isolated crossing of both $\Gamma_{2}$ (due to the nondegeneracy of the crossing form $\mathfrak{m}_{t_{0}}$) and $\Gamma_{3}$ (due to $\mathcal{N}$ having compact resolvent).
The issue with computing $\mathfrak{c}$ directly is that $(0,1)$ is a non-regular crossing of $\Gamma_{3}$; indeed, when $\lambda_{0}=0$ the eigenfunction for $\mathcal{N}+V$ is given by $\mathbf{u}_{t_{0}}=(0,v_{t_{0}})$, hence $\mathfrak{m}_{\lambda_{0}}=\left\langle\mathbf{u}_{t_{0}},\tau\mathbf{u}_{t_{0}}\right\rangle=0$. To remedy this, we will homotope the Lagrangian path through this crossing to one with only regular crossings, exploiting the fact that we can readily compute the signature of $\mathfrak{m}_{t_{0}}$ when $\lambda=0$.
To that end, we employ Theorem 4.2. In particular, there exists a $C^{2}$ curve $t=t(\lambda)$ through the point $(\lambda,t)=(0,1)$ satisfying $\dot{t}(0)=0$ and, by assumption, $\ddot{t}(0)\neq 0$. By (5.12), this curve represents the locus of points through $(\lambda,t)=(0,1)$ such that $\Upsilon_{\lambda,t(\lambda)}\cap\mathfrak{D}\neq\{0\}$. We homotope the path according to whether $t^{\prime\prime}(0)>0$ or $t^{\prime\prime}(0)<0$, see Fig. 3.
Figure 3. Schematic of the Lagrangian path (solid black) through the top left corner crossing $(\lambda,t)=(0,1)$ when the eigenvalue curve $t(\lambda)$ (blue) satisfies (a) $t^{\prime\prime}(0)<0$ and (c) $t^{\prime\prime}(0)>0$, and the new path (dashed black) to which we homotope the original path when (b) $t^{\prime\prime}(0)<0$ and (d) $t^{\prime\prime}(0)>0$. The new path has only regular crossings.
When $t^{\prime\prime}(0)<0$, we homotope the path to one having one regular crossing in the $t$ direction when $\lambda=0$, see Figs. 3(a) and 3(b). (Note that the homotoped path is indeed well-defined for $t>1$, see Remark 1.4.) In this case, using the expression for the crossing form $\mathfrak{m}_{t_{0}}$ from (5.8), since $\mathbf{u}_{t_{0}}=(0,v_{t_{0}})$ we have $n_{+}(\mathfrak{m}_{t_{0}})=1$. In this case it follows that $\mathfrak{c}=1$. On the other hand, when $t^{\prime\prime}(0)>0$ we homotope the path to one having no crossings at all, see Figs. 3(a) and 3(b). It follows that $\mathfrak{c}=0$. In summary, we have
$\mathfrak{c}=\begin{cases}+1&t^{\prime\prime}(0)<0,\\ 0&t^{\prime\prime}(0)>0.\end{cases}$ (5.23)
For the crossings along $\Gamma_{3}$ (excluding the crossing at $\lambda=0$), the crossing form (5.9) is not signed, and hence the Maslov index does not count the number of crossings. Nonetheless, the absolute value provides a lower bound for the number of crossings, i.e.
$n_{+}(\mathcal{N})\geq\big|\operatorname{Mas}(\Upsilon_{\lambda,1},\mathcal{D};\lambda\in[\varepsilon,\lambda_{\infty}])\big|.$ (5.24)
For the crossings along $\Gamma_{2}$, suppose that $t_{0}\in(\varepsilon_{0},1-\varepsilon]$ is a crossing with $\mathbf{u}_{t_{0}}\in\ker(\mathcal{N}_{t_{0}}+V_{t_{0}})$. Observe that when $\lambda=0$ the eigenvalue equations decouple into two independent equations:
$(\mathcal{N}+V_{t_{0}})\mathbf{u}=0\quad\iff\quad\begin{cases*}(\mathcal{A}+F_{t_{0}})v_{t_{0}}=0,\\ (\mathcal{A}+G_{t_{0}})u_{t_{0}}=0.\end{cases*}$ (5.25)
[…]
Summing the signatures over all crossings $t_{0}\in[\varepsilon_{0},1-\varepsilon]\subset\Gamma_{2}$ (excluding the corner crossing at $(\lambda,t)=(0,1)$), and recalling the definitions of $p_{c}$ and $q_{c}$ in (1.16), it follows that
$\operatorname{Mas}(\Upsilon_{0,t},\mathcal{D};t\in[\varepsilon_{0},1-\varepsilon])$ $=+\sum_{t_{0}}\dim\ker(\mathcal{A}+F_{t_{0}})-\sum_{t_{0}}\dim\ker(\mathcal{A}+G_{t_{0}})=q_{c}-p_{c}.$ (5.27)
Now collecting (5.24) and (5.21) together, we obtain
$n_{+}(\mathcal{N})\geq\big|\operatorname{Mas}(\Upsilon_{0,t},\mathcal{D};t\in[\varepsilon_{0},1-\varepsilon])+\mathfrak{c}\big|,$ (5.28)
and using (5.27) and (5.23), we arrive at the inequality in Theorem 1.2. ∎
Conflict of interest. The authors declare no conflict of interest.
Data availability. The data that support the findings of this study are available from the corresponding author upon reasonable request.